### v1

Abstract
Recent progress in large language models (LLMs) has shown promise in formal theorem proving, yet existing benchmarks remain limited to isolated, static proof tasks, failing to capture the iterative, engineering-intensive workflows of real-world formal mathematics libraries. Motivated by analogous advances in software engineering, we introduce the paradigm of Automated Proof Engineering (APE), which aims to automate proof engineering tasks such as feature addition, proof refactoring, and bug fixing using LLMs. To facilitate research in this direction, we present APE-Bench I, the first realistic benchmark built from real-world commit histories of Mathlib4, featuring diverse file-level tasks described in natural language and verified via a hybrid approach combining the Lean compiler and LLM-as-a-Judge. We further develop Eleanstic, a scalable parallel verification infrastructure optimized for proof checking across multiple versions of Mathlib. Empirical results on state-of-the-art LLMs demonstrate strong performance on localized edits but substantial degradation on handling complex proof engineering. This work lays the foundation for developing agentic workflows in proof engineering, with future benchmarks targeting multi-file coordination, project-scale verification, and autonomous agents capable of planning, editing, and repairing formal libraries.
[…]
Huajian Xin at
1 Introduction
Large language models (LLMs) have recently demonstrated impressive capabilities in formal mathematics, achieving state-of-the-art results on isolated theorem-proving tasks. Benchmarks such as miniF2F [17] and competition-grade problems like those in the IMO [5] showcase LLMs’ ability to reason symbolically and generate formally valid proofs. However, these benchmarks operate in static, goal-driven settings and fail to reflect the full complexity of real-world formal mathematics workflows. In practice, maintaining large formal libraries such as Mathlib4 entails more than solving standalone problems—it involves iteratively refactoring, extending, and debugging a highly interconnected and evolving codebase. This gap between isolated problem-solving and real-world proof engineering highlights the need for a fundamentally different evaluation paradigm.
While analogous limitations in code generation have led to increasingly realistic software engineering benchmarks—e.g., from HumanEval [4] to SWE-bench [10]—the formal mathematics community has yet to adopt a similar shift. In software engineering, SWE-bench reframes model evaluation as resolving realistic GitHub issues grounded in historical code evolution. We argue that formal mathematics must undergo a comparable transformation: from static theorem proving to full-fledged proof engineering, where modifications are driven by long-term development and maintenance goals rather than isolated logical challenges. Contributors to large formal libraries such as Mathlib4 engage in continuous restructuring, extension, and repair across an evolving codebase. To model this process in a way that supports LLM-based evaluation, we introduce the paradigm of Automated Proof Engineering (APE)—a setting where each task presents a natural language instruction along with a pre-edit version of Lean files, and the model is expected to generate a corresponding code patch. While human edits are not guided by explicit instructions, we distill their underlying intent into instruction–patch pairs, enabling modular, interpretable, and scalable evaluation units. Indeed, the name APE intentionally alludes to this mimicry: we aim to “ape” the structure of human proof engineering behavior in a form suitable for language models.
To operationalize this paradigm, we introduce the APE-Bench series—a staged suite of benchmarks designed to evaluate increasingly sophisticated proof engineering capabilities. As shown in Figure 1, this series progresses through three stages. APE-Bench I focuses on single-file instruction-guided edits with localized structural scope. APE-Bench II will introduce multi-file coordination and project-level verification, enabling evaluation of consistency-aware editing across interdependent modules. Finally, APE-Bench III aims to support autonomous workflows, where agents plan, edit, and iteratively repair formal libraries via feedback-driven strategies. This roadmap is not merely a scaling of difficulty, but a reflection of the layered competencies required for realistic formal proof maintenance—from local patch generation to cross-file planning and agentic self-correction.
Figure 1: The APE-Bench series: A staged roadmap for automated proof engineering. The benchmark suite progresses from single-file edits (I) to multi-file coordination (II) and ultimately to autonomous, feedback-driven workflows (III). This staged design reflects the increasing structural and semantic complexity of real-world proof engineering tasks. Dashed arrows denote prospective support for feedback-driven self-correction.
Figure 2: APE-Bench I: Benchmark structure and evaluation pipeline. (a) Unlike earlier benchmarks such as miniF2F and miniCTX, which focus on isolated theorem-proving in minimal contexts, APE-Bench I targets realistic file-level proof engineering tasks sourced from real-world Mathlib commits. These tasks involve rich, interleaved Lean declarations and demand long-range reasoning and structural precision. (b) The data pipeline begins with commit mining, diff extraction, and instruction synthesis, followed by a one-time build phase where each Mathlib version is compiled and stored using Eleanstic—a multi-version, content-addressable infrastructure for efficient Lean verification. Each commit is snapshotted as a mapping from file paths to content hashes, with all unique content stored in a global deduplicated store. During evaluation, model-generated patches are applied in isolated worktrees and verified via a two-stage process: syntactic checking with the Lean compiler and semantic checking using an LLM-as-a-Judge. This setup supports scalable, parallel, and version-aware evaluation of proof engineering capabilities.
This paper presents APE-Bench I, the first step in this roadmap. Built from thousands of real-world commits in Mathlib4, it targets localized, instruction-driven edits at the file level, providing a realistic and controlled setting for evaluating proof engineering capabilities. Its design is guided by four core goals:
  • Realism: Reflect real development practices in large-scale formal libraries;
  • Stratification: Support diverse task types and structural complexities to reveal capability gaps and preserve headroom;
  • Scalable Syntax-Semantics Evaluation: Enable syntactically efficient and semantically faithful assessment of candidate edits;
  • Continual Evolvability: Allow automated updates in step with the evolution of the underlying library.
To realize realism, we extract file-level edits from real commits in Mathlib4, preserving their structural and contextual integrity. Each edit is paired with a natural language instruction synthesized from the diff and validated through a reverse LLM-based semantic check to ensure fidelity to the original developer intent.
To ensure stratification, we construct a high-quality and diverse task set through a multi-stage filtering pipeline that removes trivial or structurally incoherent edits. To expose and verify the structural diversity of the dataset, we automatically label each task by both functional category (e.g., feature, refactor, bug fix) and difficulty level, using LLMs. These labels are empirically validated: model performance across types and difficulty bands confirms their effectiveness in revealing capability differences and guiding fine-grained evaluation.
To support scalable and semantically grounded evaluation, we adopt a two-stage protocol: syntactic checking via the Lean compiler in isolated environments (enabled by our Eleanstic infrastructure), and semantic verification via an LLM-as-a-Judge system assessing whether the patch fulfills the instruction. This approach balances robustness, automation, and evaluation cost.
Finally, to ensure continual evolvability, the entire benchmark construction pipeline—from commit mining to task synthesis, labeling, and verification—is fully automated and publicly released. It can be re-executed over future Mathlib4 versions, enabling the benchmark to remain current without manual intervention.
APE-Bench I enables a comprehensive evaluation of leading language models on realistic proof engineering tasks, revealing clear performance gaps and structural limitations. o3-mini leads by a wide margin, achieving nearly twice the semantic success rate of the next best model. While most models degrade sharply as task size increases, o3-mini remains effective even on large edits. Many models frequently produce syntactically valid but semantically incorrect patches, failing to satisfy the intended instruction. In contrast, Claude Sonnet 3.7 (thinking) achieves high semantic precision but with limited coverage. To guide future research, we categorize common semantic failure types and provide case studies that highlight recurring issues in structural planning and task interpretation.
We summarize our core contributions as follows:
  1. 1.
    We introduce the paradigm of Automated Proof Engineering (APE), which reframes formal mathematics as instruction-driven code editing aligned with real-world development and maintenance workflows.
  2. 2.
    We construct APE-Bench I, the first large-scale, file-level benchmark for proof engineering, built from over 10,000 real-world commits in Mathlib4. Tasks are automatically synthesized by mining commit diffs and generating natural-language instructions, then labeled and validated to reflect realistic intent and structural difficulty.
  3. 3.
    We develop a two-stage evaluation protocol that combines syntax-level checking via Lean compilation with semantic-level verification using LLMs. To support this, we introduce Eleanstic, a scalable, version-aware Lean infrastructure that enables efficient, stateless patch validation across thousands of commits.
  4. 4.
    We conduct a comprehensive empirical study on leading LLMs, revealing substantial performance differences, sharp degradation on structurally complex tasks, and frequent semantic failures. We further provide a taxonomy of error types and case studies to guide future research.
  5. 5.
    We will release all data, code, and infrastructure to support reproducible evaluation and continual benchmark expansion as Mathlib4 evolves.
2 Background and Related Work
2.1 Formal Proof Engineering: Practice and Challenges
Formal proof engineering refers to the disciplined application of software engineering principles to the construction, maintenance, and evolution of large-scale formal mathematical libraries [14]. In contrast to traditional theorem proving, which typically centers on isolated logical goals, proof engineering entails a broader and more structurally intricate set of tasks—such as extending libraries with new features, restructuring existing proofs (refactoring), and correcting erroneous logic (bug fixing).
Real-world formal systems, including Lean’s Mathlib [13], Isabelle/HOL’s Archive of Formal Proofs, and the seL4 microkernel verification project [11], exemplify the scale and complexity of this engineering effort. These libraries often contain thousands of interdependent definitions, theorems, and proofs, where modifying a single element may trigger cascading changes across the codebase. Ensuring logical consistency under such conditions demands significant manual labor and deep structural awareness.
While certain tools—such as iCoq for regression-proving [3]—provide partial automation, they are largely heuristic and require intensive human oversight. To date, there remains a lack of automated systems capable of performing high-fidelity edits in complex proof environments.
2.2 Machine Learning for Theorem Proving
Early machine learning efforts in formal reasoning primarily focused on tactic prediction, where proof construction is modeled as a sequential decision-making process. Systems such as HOList [2] and GamePad [8] trained neural models to select proof tactics from localized proof contexts. However, these benchmarks rely on tasks extracted from existing libraries, often involving subgoals that lack standalone mathematical significance. Consequently, they offer only limited insight into a model’s genuine reasoning capabilities.
Recent research has shifted toward end-to-end theorem proving, where models are evaluated on their ability to solve self-contained problems without intermediate scaffolding. Datasets like miniF2F [17] and FIMO [12] reflect this transition and represent a meaningful step toward assessing mathematical competence.
2.3 Limitations of Existing Benchmarks
Nevertheless, most existing benchmarks still emphasize single-goal theorem proving within static and minimal contexts, a setting that diverges significantly from the demands of real-world proof maintenance. These benchmarks evaluate success based on isolated proof completion, neglecting the broader engineering workflows needed to evolve and maintain large formal mathematics libraries.
Table 1 summarizes the conceptual differences between conventional theorem-proving benchmarks and practical proof engineering settings. In contrast to the narrow and static focus of existing benchmarks, real-world workflows involve continuous updates, structural consistency, and diverse goals such as refactoring, bug fixing, and extending functionality.
Aspect Current Benchmarks Real-world Practice
Scope Isolated, static theorem proving Dynamic, evolving large libraries
Objectives Single-attempt theorem completion Iterative refinement, multiple engineering goals
Activities Individual proof completion Feature addition, proof refactoring, bug fixing
Table 1: Comparison between conventional theorem-proving benchmarks and real-world proof engineering practices.
Several recent efforts have explored richer proof generation scenarios, such as incorporating intermediate steps, leveraging auxiliary definitions [6, 9, 15, 16, 7], or modifying existing proofs through refactoring [18] and improvement techniques [1]. However, these approaches are still typically used as tools for data augmentation or pipeline enhancement, rather than serving as rigorous evaluation frameworks. As a result, a clear gap remains in the form of systematic, task-diverse benchmarks that reflect the full scope of real-world proof engineering.
3 APE-Bench I: A Realistic Benchmark for File-Level Proof Engineering
To concretely realize the paradigm of Automated Proof Engineering (APE), we introduce APE-Bench I, a realistic and instruction-driven benchmark constructed from real-world edits in the Mathlib4 formal mathematics library. This section describes the design and construction of the benchmark, including its task format, the methodology for extracting and filtering file-level edits from version history, and the automatic synthesis of natural-language instructions and semantic labels. Together, these components enable structured, modular, and semantically faithful evaluation of language models on localized proof engineering tasks.
3.1 Task Format and Objective
Each task in APE-Bench I defines a localized file-level proof engineering problem, grounded in an actual edit to a Lean source file. The objective is to apply a valid patch that fulfills a natural-language instruction by transforming a given pre-edit file into a structurally correct and semantically meaningful post-edit version.
[…]
  • Instruction is a concise imperative-style natural language command describing the intended modification. It typically requests the addition of a definition, the restructuring of an existing proof, or the correction of a specific error.
  • PreFile contains the complete Lean source code of the target file before the edit. It provides the context necessary to interpret the instruction and generate a corresponding structural change.
  • Patch is a unified diff that encodes the intended edit, specifying insertions and deletions relative to PreFile. The diff format supports modular and interpretable editing, and facilitates automated application, verification, and evaluation.
This triplet formulation reflects the real-world dynamics of proof maintenance, where developers operate on full files with evolving structures rather than isolated subgoals. It also aligns naturally with instruction-following language models, enabling instruction-conditioned generation and evaluation of file-level structural changes. By grounding each task in a concrete diff from version history, APE-Bench I preserves the realism, coherence, and modularity essential for evaluating automated proof engineering capabilities.
3.2 Task Extraction from Mathlib Commits
(a) Token distribution by benchmark
(b) Candidate reduction across filtering stages
Figure 3: APE-Bench I: Scale and complexity of extracted tasks. (a) Compared to prior datasets, APE-Bench I tasks span significantly longer input–output contexts, reflecting greater structural and semantic complexity. (b) Starting from 19k commits, the extraction pipeline produces over 185k candidates, which are refined through multiple filtering stages to yield 10,928 high-quality tasks.
To construct realistic and structurally diverse proof engineering tasks, we extract file-level edits from the version history of Mathlib4, Lean’s primary formal mathematics library. The resulting benchmark tasks are derived from actual developer activity, ensuring contextual fidelity and development relevance. This section describes the extraction pipeline, which proceeds in four stages: commit sourcing and decomposition, chunk-level expansion, multi-stage filtering, and syntactic validation.
Step 1: Commit Sourcing and Diff Decomposition.
We begin with 19,867 Git commits made between August 2023 and March 2025, a period selected to capture stable post-migration development in Mathlib4. Each commit is decomposed into file-level diffs, isolating modified files and their corresponding changes. These diffs are further segmented into semantically scoped chunks: for modified files, we use hunk-level boundaries; for newly added files, we apply heuristics based on Lean declaration structure to determine logical segments.
Step 2: Granular Chunk Expansion.
To improve modeling tractability and enhance semantic interpretability, we decompose large diffs into smaller, contiguous edit fragments. Each fragment is applied incrementally to the evolving file state, yielding fine-grained and contextually grounded task units. This expansion increases the total number of edit candidates by nearly an order of magnitude, producing over 187,000 candidate units from the original commit set.
Step 3: Multi-Stage Filtering.
To isolate high-quality, semantically meaningful tasks, we apply a sequence of automated filters targeting three dimensions:
  • Semantic relevance: We retain only commits with recognizable editing intent, identified via conventional prefixes such as feat, fix, or refactor. Commits involving file deletions, mechanical formatting, or edits outside the Mathlib/ directory are excluded.
  • Content quality: Edits that only affect comments, import ordering, or whitespace are discarded, as they lack structural substance.
  • Structural integrity: We restrict each task to at most 100 lines of non-comment changes and a maximum input length of 16,000 tokens (measured using the GPT-4o tokenizer). Tasks must be localized, coherent, and semantically self-contained.
The most substantial reduction occurs after chunking, where structurally trivial diffs are eliminated. Subsequent filters further refine the corpus. The overall retention trajectory is visualized in Figure 3(b), illustrating each pruning stage in the pipeline.
Step 4: Compilation Validation.
To ensure structural soundness, all retained patches are compiled under their post-edit environments using Lean. Only those patches that successfully typecheck in the restored version context are included in the final benchmark. This step guarantees that each task is grounded in a valid and verifiable edit to a Lean file.
The final benchmark consists of 10,928 structurally coherent file-level tasks. As shown in Figure 3(a), these tasks exhibit significantly longer input and output sequences than prior benchmarks such as miniF2F and miniCTX. This reflects the higher structural complexity of proof engineering tasks and highlights the need for long-context reasoning and precise structural planning.
Each stage of the pipeline incrementally refines the dataset to ensure structural integrity, semantic depth, and fidelity to real-world development practices. The resulting benchmark comprises 10,928 structurally coherent file-level tasks that capture a diverse set of realistic, non-trivial modifications authored by human contributors. This dataset serves as the foundation for instruction synthesis and stratified evaluation in the subsequent sections. As shown in Figure 3(a), these tasks exhibit significantly longer input and output sequences than prior benchmarks such as miniF2F and miniCTX, underscoring the greater structural complexity of proof engineering and the need for long-context reasoning and precise structural planning.
3.3 Instruction Synthesis and Task Labelling
To enable instruction-driven modeling and facilitate stratified evaluation, each task in APE-Bench I is paired with a natural language instruction and annotated with semantic labels describing its functional category and structural difficulty. Both instruction synthesis and task labelling are performed automatically using LLMs, with verification procedures designed to ensure fidelity, clarity, and modeling relevance.
Instruction Generation.
For each filtered task derived from a Lean file diff (as described in Section 3.2), we synthesize an accompanying instruction that concisely captures the developer’s intent. We use Claude Sonnet 3.7 (thinking mode) to generate instructions from the raw diff, prompting the model to extract and summarize the intended edit in an imperative, concise, and actionable style. Instructions typically reflect changes such as introducing a new definition, modifying a proof, or restructuring an existing theorem.
To ensure semantic alignment between instruction and patch, we apply a reverse validation procedure identical to the semantic evaluation process described in Section 4.2. Specifically, the gold patch is re-evaluated against the synthesized instruction using the LLM-as-a-Judge framework. Only tasks for which the original edit satisfies the instruction are retained in the benchmark. This backward verification ensures that instructions are both meaningful and faithfully representative of the structural change.
Task Labelling.
Figure 4: Distribution of code change magnitudes across semantic labels within the test set. (a) Based on non-comment line counts from gold patches in the test set, grouped by difficulty. (b) Patch sizes grouped by functional category, based on gold commit diffs.
Figure 5: Object-level statistics of Lean patches in the test set. Distribution of added and changed def and theorem declarations, measured by count and total lines.
To enable stratified evaluation and reveal model capability differences across structurally diverse tasks, each instance in the test set is automatically annotated with semantic labels describing its functional intent and structural complexity. Specifically, we assign three types of labels:
  • Task Category, capturing functional purpose, is categorized as Feature, Refactor, or Bug Fix, aligned with common proof engineering operations.
  • Difficulty Level, estimating structural and reasoning complexity, is stratified as Very Easy, Easy, Medium, Hard, or Very Hard.
  • Task Nature, a binary label indicating whether a patch is Substantial or Superficial, used for filtering; only Substantial tasks are retained in the benchmark.
In addition to supporting dataset annotation, the labels also guide task selection: instances marked as Very Easy or Superficial are filtered out, ensuring that the benchmark retains only structurally substantive and semantically meaningful edits. Beyond this filtering role, the labels enable stratified evaluation across both functional categories and complexity levels, forming the foundation for the fine-grained performance analysis in Section 5.3. We next assess the effectiveness of these labels by analyzing their correlation with code magnitude and declaration-level structure.
Statistical Validation of Difficulty Labels.
Figure 4(a) shows that while Easy tasks are generally small, Medium and Hard tasks span a broad range of edit sizes, with no strict separation based on line count. This indicates that difficulty labels capture more than superficial metrics—they reflect underlying semantic and structural complexity. This is further confirmed by object-level patterns in Figure 5. Medium and Hard tasks involve more declarations and larger edits per object, particularly for theorems, suggesting greater logical coordination and proof effort. These trends align with empirical results in Section 5.3, where model performance declines sharply from Easy to Hard, validating the semantic significance of the difficulty labels.
Functional Differentiation by Task Category.
Task categories also reveal distinct structural roles. As shown in Figure 4(b), Feature tasks are larger and often involve new declarations, while Refactor and Bug Fix edits are smaller and focus on modifying existing code. Object-level statistics support this: Feature tasks add theorems and definitions, while the other two primarily adjust existing defs. These differences reflect practical development workflows. The small scale of Refactor and Bug Fix edits stems in part from filtering heuristics, which exclude trivial or mechanical changes. As such, retained tasks across all categories represent meaningful engineering actions.
4 Evaluation Infrastructure and Protocol
Figure 6: Evaluation pipeline for APE-Bench I. A model-generated patch undergoes syntactic verification via the Lean compiler and semantic judgement via an LLM-as-a-Judge. A task is marked as successful only if both stages are passed.
To enable reliable and automated assessment of model-generated edits in APE-Bench I, we establish a two-stage evaluation framework that reflects the dual demands of proof engineering: syntactic correctness and semantic adequacy. Each candidate patch is evaluated for both grammatical validity under Lean’s type system and alignment with the task’s instruction-level intent.
The first stage, syntactic verification, checks whether the resulting file of the generated pass the Lean verification within the corresponding versioned environment. This process is implemented via Eleanstic, a scalable verification infrastructure that supports reproducible, parallelizable, and version-aware Lean verification infrastructure (Section 4.1).
The second stage, semantic judgement, determines whether the patch satisfies the instruction as intended. This is carried out by a large language model acting as an instruction-conditioned judge, which reasons over the pre-edit file, the patch, and the goal description to determine task success (Section 4.2).
Figure 6 presents the overall evaluation pipeline. A task is marked as successful only if a model-generated patch passes both verification stages. This protocol provides a standardized, interpretable, and scalable basis for computing success metrics, and enables detailed analysis of model performance across syntax-level and semantics-level axes.
4.1 Syntactic Verification via Eleanstic
Syntactic verification in APE-Bench I requires compiling model-generated patches within the exact versioned environment from which each task was derived. In large-scale formal libraries like Mathlib4, where definitions and dependencies evolve rapidly across commits, such version fidelity is critical to ensure meaningful evaluation. However, this fidelity introduces significant engineering complexity: naively rebuilding the library for every historical commit results in both prohibitive computational cost and unsustainable storage overhead. A single clean build of Mathlib4 can take several hours to complete and produce multiple gigabytes of intermediate artifacts. When scaled to thousands of commits—each potentially associated with thousands of patch candidates and evaluated across multiple models—this cost compounds rapidly, impeding both throughput and reproducibility.
To address this, we introduce Eleanstic, a dedicated infrastructure for scalable, version-aware syntactic verification. Eleanstic decouples the high-cost, stateful process of building Mathlib from the low-cost, stateless process of validating edits. It does so by collapsing each post-build state into a compact, content-addressable snapshot that can later be restored to yield a faithful Lean environment for verification. This architecture significantly reduces redundant computation, minimizes on-disk storage usage, and supports efficient parallel evaluation across thousands of distinct versions.
Snapshot Generation.
For each unique commit, Eleanstic first performs a clean lake build and recursively traverses the resulting worktree. Every file is hashed using SHA-256 and stored in a content-addressable storage (CAS) structure with a two-level hash-based directory layout. In parallel, a binary snapshot is created to record the mapping from relative file paths to content hashes and file types (regular or symbolic). This snapshot, encoded using fixed-width binary records, serves as a complete and portable blueprint of the post-build environment, ensuring reproducibility while minimizing metadata overhead.
Snapshot Restoration.
At verification time, the appropriate snapshot is loaded and its contents restored into a temporary, isolated worktree that exactly replicates the post-build file structure of the original compilation. Files are fetched from the global CAS and written to disk. This process is stateless and fully parallelizable, enabling fast switching across historical commits without requiring repeated builds. Empirically, Eleanstic restores a full Lean environment in under one second on SSD-backed systems. To support long-term scalability, Eleanstic performs aggressive global deduplication across all snapshots. Across 3734 commits spanning August 2023 to March 2025, it compresses 15.6 TB of build artifacts into just 1.1 TB, yielding a storage reduction of over 93%. This compression enables persistent caching of every historical state required for evaluation, without exhausting disk resources.
Patch Verification.
After restoring the environment, the model-generated patch (see Section 5.1) is applied to the target Lean file, and the resulting content is saved to a temporary file. Verification follows a strict file-level isolation policy: rooted in the restored Mathlib worktree, the Lean compiler compiles only the temporary file without affecting any files in the restored environment. Additionally, any temporary file that attempts to import the target file is discarded to prevent unintended leakage across file boundaries. A patch is considered syntactically valid if it compiles successfully within the corresponding version of Mathlib4 and does not emit warnings.
4.2 Semantic Judgement via LLM-as-a-Judge
While syntactic verification ensures that a model-generated patch is type-correct under the Lean compiler, it does not guarantee that the patch fulfills the intended transformation described in the natural language instruction. Common failure modes include omitting key components, misinterpreting the instruction, or modifying unrelated content. To address this gap, we introduce a semantic evaluation protocol based on the LLM-as-a-Judge paradigm, which assesses whether a patch correctly implements the specified task.
Model and Prompt.
We use Claude Sonnet 3.7 (thinking mode) as the semantic evaluator. Given a task triplet—comprising the instruction, the pre-edit Lean file, and the candidate patch—the model is prompted to assess whether the patch fulfills the instruction in both structure and intent. The prompt encourages multi-step reasoning over the instruction, analysis of the proposed changes, and consistency verification. To ensure output stability and reduce overfitting to prompt phrasing, the full evaluation prompt is designed to be modular and instruction-agnostic.
Voting Strategy and Decision Criteria.
To improve robustness, we adopt a sample@4 strategy: four independent judgments are sampled from Claude for each evaluation instance. The final verdict is determined by majority vote, where a task is considered semantically successful only if the majority of samples agree that all subcomponents of the instruction are correctly implemented. This approach balances reliability and scalability while maintaining consistency across model outputs. All semantic evaluation results reported in this work are based on this protocol.
5 Experiments
5.1 Experimental Setup
We evaluate seven general-purpose language models on APE-Bench I using a standardized evaluation pipeline. The task for each model is to generate a Lean patch that fulfills a natural language instruction, based on a complete pre-edit source file. Success is measured via a two-stage evaluation combining syntactic verification and semantic judgement.
Models.
We evaluate a range of state-of-the-art, general-purpose LLMs, selected to represent both long-context reasoning and standard generation capabilities:
  • Reasoning models: OpenAI o3-mini, Claude Sonnet 3.7 (thinking), DeepSeek R1.
  • Non-reasoning models: Claude Sonnet 3.7, DeepSeek V3 (250324), GPT-4o, Doubao 1.5 Pro.
All models are zero-shot, without any task-specific fine-tuning for proof engineering.
Patch Normalization via DiffRepair.
Model outputs are produced in unified diff format, which often contains structurally imprecise or misaligned edits (e.g., incorrect line anchors, out-of-date contexts). To ensure reliable application and verification, we introduce a post-processing phase called DiffRepair, which rewrites imprecise diffs into structurally consistent patches. This mechanism operates in two phases: first, it localizes the intended edit position using fuzzy anchor matching; then it reconstructs the patch with preserved structural and contextual integrity. All generated patches are normalized through DiffRepair before subsequent compilation or evaluation. (See Appendix A for full technical details.)
Sampling Methodology.
We adopt the standard pass@$k$ evaluation [4] to estimate task success rates. For each task, $n=20$ candidate patches are sampled, and pass@16 is computed using the unbiased estimator:
[…]
where $c\leq n$ is the number of patches passing both syntactic and semantic verification.
Temperature Settings.
For models supporting temperature control, we evaluate outputs under two sampling temperatures, $T\in\{0.0,0.6\}$. Empirical results indicate that $T=0.6$ consistently improves output diversity and task coverage. For reasoning models that do not expose temperature parameters, we adopt default or recommended configurations: o3-mini is run with medium reasoning effort, while Claude Sonnet (thinking) is allocated 6,000 tokens of thinking budget to enable extended context comprehension.
5.2 Main Results
We evaluate model performance using the pass@16 metric, which estimates the probability that at least one of 16 generated candidates satisfies specified criteria. Specifically, we distinguish between:
  • Lean Compile pass@16: The proportion of tasks for which at least one candidate compiles successfully using Lean 4, validating syntactic and structural soundness.
  • LLM-Judge pass@16: The proportion of tasks for which at least one compiled candidate additionally satisfies the instruction according to semantic evaluation via LLM-as-a-Judge.
  • Semantic Drop: The percentage decrease from Lean Compile to LLM-Judge success, reflecting how often syntactically valid outputs fail to meet the intended edits.
Table 2 summarizes the performance of seven general-purpose models. Among all candidates, o3-mini achieves the highest overall success rate, with 33.58% of tasks completed successfully under full evaluation. It also shows the highest syntactic pass rate at 48.38%, indicating its ability to generate compilable edits across a broad range of task types. However, its semantic drop-the proportion of compilable patches that fail to meet instruction-level requirements-is also substantial at over 30%, suggesting frequent deviation from intended modifications.
In contrast, Claude Sonnet 3.7 (thinking) exhibits perfect semantic consistency with a semantic drop of zero, yet achieves a significantly lower overall success rate due to reduced syntactic coverage. GPT-4o and Doubao 1.5 Pro show the opposite trend: both suffer from large semantic drops of 49.31% and 32.21%, respectively. These patterns highlight fundamental differences in model behavior—while some models excel at generating well-formed Lean code, they often fail to capture the semantic intent of the editing instruction.
Model Lean Compile pass@16 LLM-Judge pass@16 Semantic Drop (%)
o3-mini 48.38 33.58 30.58
Claude Sonnet 3.7 (thinking) 17.73 17.73 0.00
DeepSeek R1 17.67 15.85 10.30
DeepSeek V3 16.23 15.72 3.11
Claude Sonnet 3.7 15.85 14.85 6.31
GPT-4o 18.94 9.60 49.31
Doubao 1.5 Pro 9.70 6.58 32.21
Table 2: Overall model performance on APE-Bench I using pass@16. Semantic Drop is computed as the percentage of Lean-compilable outputs that fail LLM-based semantic validation.
Figure 7 further illustrates model behavior across different pass@$k$ thresholds. The left panel reports syntactic success (Lean Compile), while the right panel shows semantic success (LLM-Judge) conditioned on compilation. o3-mini consistently outperforms other models across $k$ values in both dimensions, confirming its broad sampling diversity and task coverage. Claude Sonnet 3.7 (thinking) remains flat across $k$, reflecting its low-variance generation strategy.
Figure 7: Model performance across pass@$k$ values. Left: Lean Compile pass@$k$ (syntactic validity); Right: LLM Judge pass@$k$ (semantic correctness conditioned on compilation). The gap between the two reflects semantic degradation.
Notably, even the best-performing model—o3-mini—succeeds on only about one-third of the benchmark tasks, despite the fact that all tasks are constrained to single-file edits of no more than 100 lines. This highlights the substantial gap that remains between present capabilities and the demands of real-world proof engineering. In the following sections, we analyze this failure spectrum in greater detail by examining the structural and semantic robustness of model outputs.
5.3 Structural Robustness and Task-Scale Sensitivity
While overall success rates reveal substantial variation across models, they do not capture how model performance scales with the structural complexity of the underlying tasks. In this section, we analyze the extent to which models can handle edits of increasing size and structural scope, assess their ability to produce appropriately scaled patches, and evaluate their robustness across task types, difficulty levels, and mathematical domains.
Performance Degradation on Larger Edits.
We first group benchmark tasks by the size of their ground-truth patches and evaluate model pass@16 within each bin. As shown in Figure 8, most models exhibit sharp performance degradation as the size of the target edit increases. While all models achieve moderate to high success rates on small-scale edits (1–5 lines), performance drops significantly for medium (11–20 lines) and large-scale (>20 lines) tasks. In particular, o3-mini is the only model to retain non-trivial success on edits exceeding 50 lines, with a pass@16 about 25%, while all other models collapse to near-zero in this regime. These results highlight the inability of most LLMs to scale beyond localized modifications.
Figure 8: Model pass@16 success rate grouped by ground-truth patch size. Larger edits cause sharp performance degradation for most models. Only o3-mini maintains partial success on tasks with more than 20 lines of changes.
Patch Length Alignment with Ground-Truth.
Figure 9 evaluates whether model-generated patches exhibit scale-aware behavior by comparing their length to the corresponding human edits. Most models over-edit small tasks (1–5 lines), producing longer-than-needed patches. As task size increases, generated patch sizes tend to align more closely with ground truth—except in models like GPT-4o and Doubao 1.5 Pro, which increasingly under-edit larger patches. This under-generation corresponds to their low semantic success on structurally complex tasks.
In contrast, o3-mini and other strong models (Claude Sonnet, DeepSeek variants) exhibit broader variance and better coverage across patch sizes. Their outputs span both under- and over-editing, reflecting more diverse structural hypotheses. Notably, o3-mini frequently over-edits larger patches, which increases numeric deviation but improves semantic success by avoiding missing components—the dominant failure mode among conservative models. These results suggest that scale diversity, rather than strict length matching, is more critical for robust proof engineering performance.
Figure 9: Patch length alignment across models (log-log scale). Each panel plots the number of non-comment lines changed by the model (y-axis) versus the human reference patch (x-axis). The dashed diagonal indicates ideal alignment.
Stratified Performance by Task Type, Difficulty, and Domain.
To further investigate structural sensitivity, we evaluate model performance across different task categories, difficulty levels, and mathematical domains. Figure 10 presents a unified heatmap showing evaluation pass@16 for each model across these dimensions. Tasks labeled as bug fixes are handled reasonably well by all models, likely due to their localized and low-risk nature. Refactor tasks reveal stronger model differentiation: Claude variants perform relatively well due to their structural preservation bias, whereas GPT-4o and Doubao underperform. Feature addition tasks are the most challenging, requiring new definitions, theorems, and often multi-declaration coherence; only o3-mini exhibits notable success in this category.
Difficulty-based analysis further reveals a sharp capability cliff. While all models succeed on easy tasks, only o3-mini retains significant performance on medium and hard cases. Hard tasks, involving large edits or cross-declaration dependencies, remain virtually unsolved by all models except o3-mini. Finally, we observe domain-based variation: linear algebra and category theory consistently challenge most models, while tasks from algebra or data modules are comparatively tractable. These differences reflect varying structural depth and abstraction levels across formal math subfields.
Figure 10: Heatmap of evaluation pass@16 success rates across different task types, difficulty levels, and mathematical modules. Column widths represent the proportion of each class in the test set. Color intensity indicates model performance.
Summary.
Taken together, these results indicate that most current models lack the structural robustness necessary for realistic proof engineering. Their success is largely limited to small-scale tasks and structurally shallow modifications. As patch size increases or task complexity deepens, performance degrades sharply. Even when a model produces a syntactically valid edit, it often fails to match the structural scope expected of a real engineer. In the next section, we examine whether these structural failures also correspond to deeper semantic errors.
5.4 Semantic Robustness and Failure Characterization
Despite producing syntactically valid Lean code, many model-generated patches fail to fulfill the semantic requirements expressed in the instruction. In this section, we examine the nature of these semantic failures in detail. We analyze patch repairability, categorize common failure modes, and compare error distributions across models to identify systematic weaknesses in current approaches.
Structural Repairability of Model-Generated Patches.
Raw patches generated by LLMs often contain formatting inconsistencies, invalid anchors, or malformed context that prevent successful application. To assess structural robustness, we measure patch application success before and after applying two forms of automated repair: exact match and fuzzy match (See Appendix A for full technical details.). As shown in Table 3, raw apply success rates remain low for all models, ranging from 12% to 16%, indicating poor structural precision in the initial output. Exact match repair offers moderate improvement for DeepSeek models, but fails to resolve structural errors in o3-mini or GPT-4o outputs. Fuzzy match repair significantly boosts success across the board, particularly for DeepSeek R1, Doubao, and Claude Sonnet 3.7. These results highlight the necessity of structural repair mechanisms in enabling meaningful evaluation and practical usability.
Model Raw Apply Success Exact Match Repair Fuzzy Match Repair
DeepSeek V3 15.60% 52.08% 83.01%
DeepSeek R1 14.82% 49.46% 83.01%
Doubao 1.5 Pro 15.62% 13.60% 81.64%
Claude Sonnet 3.7 13.39% 13.72% 70.30%
Claude Sonnet 3.7 (thinking) 12.69% 14.22% 69.19%
o3-mini 14.22% 22.34% 58.78%
GPT-4o 12.53% 6.58% 42.95%
Table 3: Patch success rates before and after structural repair. Exact match applies conservative anchor correction; fuzzy match enables tolerant application using semantic alignment.
Semantic Error Taxonomy and Cross-Model Distribution.
To understand why many syntactically valid patches nonetheless fail semantically, we analyze failure cases using a taxonomy of five high-level semantic error types (Table 4). This categorization enables structured evaluation across models and highlights recurring patterns of semantic fragility.
Error Class Description
Component Omission Required definitions, proofs, or examples are missing
Incompleteness Issue The edit is structurally partial, unfinished, or placeholder
Mathematical Error The logic is invalid or inconsistent with Lean’s formal semantics
Quality Defect Redundant or unabstracted edits, or lack of documentation
Requirement Mismatch Misalignment with the intended instruction
Table 4: Semantic error classes and their typical manifestations.
Figure 11: Distribution of semantic error types across models (log-scaled y-axis). Component omission is the most common error overall. GPT-4o and o3-mini frequently produce incomplete or mathematically incorrect edits, while Claude Sonnet 3.7 (thinking) exhibits the lowest error incidence.
Figure 11 presents the frequency of each error type as a proportion of failure cases per model. The most pervasive issue is component omission, accounting for a significant share of failures across all models. This indicates a widespread tendency to under-produce or neglect structurally necessary elements—such as required definitions, lemmas, or proofs—even when the instruction clearly specifies them. Beyond omission, GPT-4o and o3-mini suffer from notably high rates of incompleteness and mathematical errors. These failures suggest limited planning ability when tasked with multi-step edits, and difficulty in maintaining formal soundness under structural pressure. By contrast, Claude Sonnet 3.7 (thinking) demonstrates a markedly lower error profile across all categories.
Summary.
Our findings indicate that semantic failures in LLM-generated patches are not random but structurally patterned. The majority stem from incomplete or underspecified edits, often due to scale underestimation or partial instruction comprehension. While fuzzy repair can address syntactic alignment, it cannot compensate for missing logic or misaligned intent. Addressing these issues will require models that plan multi-step edits, verify task satisfaction, and iteratively refine their output—an agenda we return to in our discussion of future directions.
6 Conclusion
This work introduces the paradigm of Automated Proof Engineering (APE) and presents APE-Bench I, the first large-scale benchmark grounded in real-world proof development workflows from Mathlib4. By reframing formal mathematics as instruction-driven code editing, and combining Lean compilation with LLM-based semantic evaluation, APE-Bench I offers a rigorous and scalable framework for evaluating the structural and semantic capabilities of LLMs in realistic proof engineering tasks. Our empirical results reveal significant performance gaps among state-of-the-art models, particularly under structurally complex scenarios, and underscore the need for future systems that can plan, verify, and refine edits across evolving formal libraries. This benchmark establishes a strong foundation for developing next-generation autonomous agents capable of practical and scalable formal proof maintenance.
Appendix A DiffRepair: Fault-Tolerant Patch Recovery for LLM Outputs
A.1 Motivation and Overview
Model-generated patches in unified diff format frequently fail to apply cleanly to real-world codebases due to misaligned line anchors, outdated contextual information, or excessive reliance on idealized formatting. This undermines the practical viability of LLM-generated edits in structured formal environments like Lean. To address this, we develop DiffRepair, a two-stage recovery system that transforms noisy model diffs into clean, structure-preserving, and context-valid patches that can be compiled and semantically evaluated.
DiffRepair adopts a tolerant matching and reconstruction strategy. Rather than enforcing exact line-based alignment, it reconstructs the intent behind the edit, anchors it in the correct structural region, and regenerates a diff conforming to unified format constraints. This allows it to serve as a bridge between neural code generation and downstream compilation pipelines.
A.2 Algorithmic Workflow
1. Intent Localization.
Given a noisy diff, DiffRepair first identifies where in the original Lean file the intended modification should take place. It uses a multi-granularity fuzzy matching strategy based on contextual anchors surrounding each diff hunk. Anchors are matched using:
  • Exact or prefix-based line similarity.
  • Normalized token-level matching with whitespace and casing tolerance.
  • Structural proximity, emphasizing declaration-aligned contexts (e.g., def, theorem blocks).
Matches are scored using a dynamic programming model optimizing for structural alignment and modification preservation, ensuring minimal distortion to the intended patch intent.
2. Patch Reconstruction.
Once the target region is localized, DiffRepair reconstructs a clean diff by:
  • Re-aligning added and deleted lines to structurally valid positions.
  • Augmenting missing context to satisfy unified diff constraints.
  • Resolving line number offsets and conflicting hunk overlaps.
  • Merging multiple hunks into a minimal, coherent patch that preserves original semantics.
Reconstructed patches are guaranteed to apply via standard tools such as patch or git apply, enabling seamless integration into downstream pipelines.
A.3 Design Features
DiffRepair is designed to satisfy the following desiderata:
  • Robustness to noisy generation: Handles line number inaccuracies, formatting inconsistencies, and incomplete context.
  • Structure preservation: Ensures generated diffs align with Lean’s structural boundaries and avoid cross-declaration contamination.
  • Automation-ready: Requires no human-in-the-loop intervention and supports batch repair at scale.
  • Efficiency: Performs fast multi-hunk matching and reconstruction with low overhead.

### v2

@@
 We further develop Eleanstic, a scalable parallel verification infrastructure optimized for proof checking across multiple versions of Mathlib.
-Empirical results on state-of-the-art LLMs demonstrate strong performance on localized edits but substantial degradation on handling complex proof engineering.
+Empirical results on state-of-the-art LLMs reveal a syntax–semantics gap: while some models perform well on localized edits, overall success rates drop sharply on structurally complex tasks.
 This work lays the foundation for developing agentic workflows in proof engineering, with future benchmarks targeting multi-file coordination, project-scale verification, and autonomous agents capable of planning, editing, and repairing formal libraries.
 […]
-Huajian Xin at
+[…]
 1 Introduction
 Large language models (LLMs) have recently demonstrated impressive capabilities in formal mathematics, achieving state-of-the-art results on isolated theorem-proving tasks.
-Benchmarks such as miniF2F [17] and competition-grade problems like those in the IMO [5] showcase LLMs’ ability to reason symbolically and generate formally valid proofs.
-However, these benchmarks operate in static, goal-driven settings and fail to reflect the full complexity of real-world formal mathematics workflows.
-In practice, maintaining large formal libraries such as Mathlib4 entails more than solving standalone problems—it involves iteratively refactoring, extending, and debugging a highly interconnected and evolving codebase.
+However, existing benchmarks operate in static, goal-driven settings and fail to reflect the full complexity of real-world formal mathematics workflows.
+Maintaining large formal libraries such as Mathlib4 entails more than solving standalone problems—it involves iteratively refactoring, extending, and debugging an interconnected and evolving codebase.
 This gap between isolated problem-solving and real-world proof engineering highlights the need for a fundamentally different evaluation paradigm.
-While analogous limitations in code generation have led to increasingly realistic software engineering benchmarks—e.g., from HumanEval [4] to SWE-bench [10]—the formal mathematics community has yet to adopt a similar shift.
-In software engineering, SWE-bench reframes model evaluation as resolving realistic GitHub issues grounded in historical code evolution.
-We argue that formal mathematics must undergo a comparable transformation: from static theorem proving to full-fledged proof engineering, where modifications are driven by long-term development and maintenance goals rather than isolated logical challenges.
-Contributors to large formal libraries such as Mathlib4 engage in continuous restructuring, extension, and repair across an evolving codebase.
-To model this process in a way that supports LLM-based evaluation, we introduce the paradigm of Automated Proof Engineering (APE)—a setting where each task presents a natural language instruction along with a pre-edit version of Lean files, and the model is expected to generate a corresponding code patch.
-While human edits are not guided by explicit instructions, we distill their underlying intent into instruction–patch pairs, enabling modular, interpretable, and scalable evaluation units.
-Indeed, the name APE intentionally alludes to this mimicry: we aim to “ape” the structure of human proof engineering behavior in a form suitable for language models.
-To operationalize this paradigm, we introduce the APE-Bench series—a staged suite of benchmarks designed to evaluate increasingly sophisticated proof engineering capabilities.
-As shown in Figure 1, this series progresses through three stages.
-APE-Bench I focuses on single-file instruction-guided edits with localized structural scope.
-APE-Bench II will introduce multi-file coordination and project-level verification, enabling evaluation of consistency-aware editing across interdependent modules.
+While analogous limitations in code generation have led to increasingly realistic software engineering benchmarks-e.g., from HumanEval [3] to SWE-bench [7]-the formal mathematics community has yet to adopt a similar shift.
+We introduce the paradigm of Automated Proof Engineering (APE)—a setting where each task presents a natural language instruction along with a pre-edit version of Lean files, and the model generates a corresponding code patch.
+While human edits are not guided by explicit instructions, we distill their underlying intent into instruction–patch pairs, enabling interpretable and scalable evaluation units.
+To operationalise this paradigm, we introduce the APE-Bench series—a staged suite of benchmarks for increasingly sophisticated proof engineering capabilities.
+As shown in Figure 1, APE-Bench I focuses on single-file instruction-guided edits with localised scope.
+APE-Bench II will introduce multi-file coordination and project-level verification.
 Finally, APE-Bench III aims to support autonomous workflows, where agents plan, edit, and iteratively repair formal libraries via feedback-driven strategies.
-This roadmap is not merely a scaling of difficulty, but a reflection of the layered competencies required for realistic formal proof maintenance—from local patch generation to cross-file planning and agentic self-correction.
 Figure 1: The APE-Bench series: A staged roadmap for automated proof engineering.
-The benchmark suite progresses from single-file edits (I) to multi-file coordination (II) and ultimately to autonomous, feedback-driven workflows (III).
-This staged design reflects the increasing structural and semantic complexity of real-world proof engineering tasks.
-Dashed arrows denote prospective support for feedback-driven self-correction.
+The benchmark suite progresses from single-file edits (APE-Bench I) to multi-file coordination (APE-Bench II) and ultimately to autonomous workflows (APE-Bench III).
+This staged design reflects the increasing complexity of real-world proof engineering tasks.
 Figure 2: APE-Bench I: Benchmark structure and evaluation pipeline.
-(a) Unlike earlier benchmarks such as miniF2F and miniCTX, which focus on isolated theorem-proving in minimal contexts, APE-Bench I targets realistic file-level proof engineering tasks sourced from real-world Mathlib commits.
-These tasks involve rich, interleaved Lean declarations and demand long-range reasoning and structural precision.
-(b) The data pipeline begins with commit mining, diff extraction, and instruction synthesis, followed by a one-time build phase where each Mathlib version is compiled and stored using Eleanstic—a multi-version, content-addressable infrastructure for efficient Lean verification.
-Each commit is snapshotted as a mapping from file paths to content hashes, with all unique content stored in a global deduplicated store.
-During evaluation, model-generated patches are applied in isolated worktrees and verified via a two-stage process: syntactic checking with the Lean compiler and semantic checking using an LLM-as-a-Judge.
-This setup supports scalable, parallel, and version-aware evaluation of proof engineering capabilities.
+(a) Unlike earlier benchmarks focusing on isolated theorem-proving, APE-Bench I targets realistic file-level proof engineering tasks from real-world Mathlib commits.
+(b) The data pipeline includes commit mining, diff extraction, instruction synthesis, and verification using Eleanstic-a multi-version infrastructure for efficient Lean verification.
 This paper presents APE-Bench I, the first step in this roadmap.
-Built from thousands of real-world commits in Mathlib4, it targets localized, instruction-driven edits at the file level, providing a realistic and controlled setting for evaluating proof engineering capabilities.
+Built from thousands of real-world commits in Mathlib4, it targets localised, instruction-driven edits at the file level.
 Its design is guided by four core goals:
-  • Realism: Reflect real development practices in large-scale formal libraries;
-  • Stratification: Support diverse task types and structural complexities to reveal capability gaps and preserve headroom;
-  • Scalable Syntax-Semantics Evaluation: Enable syntactically efficient and semantically faithful assessment of candidate edits;
-  • Continual Evolvability: Allow automated updates in step with the evolution of the underlying library.
-To realize realism, we extract file-level edits from real commits in Mathlib4, preserving their structural and contextual integrity.
-Each edit is paired with a natural language instruction synthesized from the diff and validated through a reverse LLM-based semantic check to ensure fidelity to the original developer intent.
-To ensure stratification, we construct a high-quality and diverse task set through a multi-stage filtering pipeline that removes trivial or structurally incoherent edits.
-To expose and verify the structural diversity of the dataset, we automatically label each task by both functional category (e.g., feature, refactor, bug fix) and difficulty level, using LLMs.
-These labels are empirically validated: model performance across types and difficulty bands confirms their effectiveness in revealing capability differences and guiding fine-grained evaluation.
-To support scalable and semantically grounded evaluation, we adopt a two-stage protocol: syntactic checking via the Lean compiler in isolated environments (enabled by our Eleanstic infrastructure), and semantic verification via an LLM-as-a-Judge system assessing whether the patch fulfills the instruction.
-This approach balances robustness, automation, and evaluation cost.
-Finally, to ensure continual evolvability, the entire benchmark construction pipeline—from commit mining to task synthesis, labeling, and verification—is fully automated and publicly released.
-It can be re-executed over future Mathlib4 versions, enabling the benchmark to remain current without manual intervention.
-APE-Bench I enables a comprehensive evaluation of leading language models on realistic proof engineering tasks, revealing clear performance gaps and structural limitations.
-o3-mini leads by a wide margin, achieving nearly twice the semantic success rate of the next best model.
-While most models degrade sharply as task size increases, o3-mini remains effective even on large edits.
-Many models frequently produce syntactically valid but semantically incorrect patches, failing to satisfy the intended instruction.
-In contrast, Claude Sonnet 3.7 (thinking) achieves high semantic precision but with limited coverage.
-To guide future research, we categorize common semantic failure types and provide case studies that highlight recurring issues in structural planning and task interpretation.
+  • Realism: Reflect real development practices by extracting file-level edits from actual Mathlib4 commits, preserving their structural integrity.
+  • Stratification: Support diverse task types and complexities through a multi-stage filtering pipeline that removes trivial edits and categorizes tasks by function and difficulty.
+  • Scalable Syntax-Semantics Evaluation: Enable efficient assessment through syntactic verification via the Lean compiler and semantic judgement via an LLM-as-a-judge system.
+  • Continual Evolvability: Support automated updates with the evolution of the underlying library through fully automated construction pipelines.
+APE-Bench I provides a rigorous framework for evaluating state-of-the-art language models on realistic proof engineering tasks.
+Experimental results show that although Gemini 2.5 Pro Preview achieves the highest semantic accuracy under LLM-based evaluation, it successfully completes only one-sixth of the benchmark tasks and fails entirely on the most challenging ones.
+Most models exhibit significant performance degradation as task complexity increases, often generating patches that are syntactically valid but semantically flawed.
+These findings underscore a substantial gap between current model capabilities and the demands of real-world proof engineering.
 We summarize our core contributions as follows:
   1. 1.
-    We introduce the paradigm of Automated Proof Engineering (APE), which reframes formal mathematics as instruction-driven code editing aligned with real-world development and maintenance workflows.
+    We present Automated Proof Engineering (APE)—a paradigm that casts LLM-based formal reasoning as instruction-driven code edits reflecting real-world development practices.
   2. 2.
     We construct APE-Bench I, the first large-scale, file-level benchmark for proof engineering, built from over 10,000 real-world commits in Mathlib4.
-Tasks are automatically synthesized by mining commit diffs and generating natural-language instructions, then labeled and validated to reflect realistic intent and structural difficulty.
   3. 3.
-    We develop a two-stage evaluation protocol that combines syntax-level checking via Lean compilation with semantic-level verification using LLMs.
-To support this, we introduce Eleanstic, a scalable, version-aware Lean infrastructure that enables efficient, stateless patch validation across thousands of commits.
+    We develop a two-stage evaluation protocol combining syntactic verification via Lean compilation with semantic judgement via an LLM-as-a-judge system.
   4. 4.
-    We conduct a comprehensive empirical study on leading LLMs, revealing substantial performance differences, sharp degradation on structurally complex tasks, and frequent semantic failures.
-We further provide a taxonomy of error types and case studies to guide future research.
+    We conduct a comprehensive empirical study on state-of-the-art LLMs, revealing substantial performance differences, sharp degradation on complex tasks, and frequent semantic failures.
   5. 5.
-    We will release all data, code, and infrastructure to support reproducible evaluation and continual benchmark expansion as Mathlib4 evolves.
+    We release all data, code, and infrastructure to support reproducible evaluation and continual benchmark expansion as Mathlib4 evolves.
 2 Background and Related Work
-2.1 Formal Proof Engineering: Practice and Challenges
 Formal proof engineering refers to the disciplined application of software engineering principles to the construction, maintenance, and evolution of large-scale formal mathematical libraries [10].
-In contrast to traditional theorem proving, which typically centers on isolated logical goals, proof engineering entails a broader and more structurally intricate set of tasks—such as extending libraries with new features, restructuring existing proofs (refactoring), and correcting erroneous logic (bug fixing).
-Real-world formal systems, including Lean’s Mathlib [13], Isabelle/HOL’s Archive of Formal Proofs, and the seL4 microkernel verification project [11], exemplify the scale and complexity of this engineering effort.
-These libraries often contain thousands of interdependent definitions, theorems, and proofs, where modifying a single element may trigger cascading changes across the codebase.
-Ensuring logical consistency under such conditions demands significant manual labor and deep structural awareness.
-While certain tools—such as iCoq for regression-proving [3]—provide partial automation, they are largely heuristic and require intensive human oversight.
-To date, there remains a lack of automated systems capable of performing high-fidelity edits in complex proof environments.
-2.2 Machine Learning for Theorem Proving
-Early machine learning efforts in formal reasoning primarily focused on tactic prediction, where proof construction is modeled as a sequential decision-making process.
-Systems such as HOList [2] and GamePad [8] trained neural models to select proof tactics from localized proof contexts.
-However, these benchmarks rely on tasks extracted from existing libraries, often involving subgoals that lack standalone mathematical significance.
-Consequently, they offer only limited insight into a model’s genuine reasoning capabilities.
-Recent research has shifted toward end-to-end theorem proving, where models are evaluated on their ability to solve self-contained problems without intermediate scaffolding.
-Datasets like miniF2F [17] and FIMO [12] reflect this transition and represent a meaningful step toward assessing mathematical competence.
-2.3 Limitations of Existing Benchmarks
-Nevertheless, most existing benchmarks still emphasize single-goal theorem proving within static and minimal contexts, a setting that diverges significantly from the demands of real-world proof maintenance.
-These benchmarks evaluate success based on isolated proof completion, neglecting the broader engineering workflows needed to evolve and maintain large formal mathematics libraries.
-Table 1 summarizes the conceptual differences between conventional theorem-proving benchmarks and practical proof engineering settings.
-In contrast to the narrow and static focus of existing benchmarks, real-world workflows involve continuous updates, structural consistency, and diverse goals such as refactoring, bug fixing, and extending functionality.
+In contrast to traditional theorem proving, which typically centres on isolated logical goals, proof engineering entails a broader and more structurally intricate set of tasks, such as extending libraries with new features, restructuring existing proofs (refactoring), and correcting erroneous logic (bug fixing).
+For a more comprehensive discussion of proof engineering practices and challenges, see Appendix .
+Most existing benchmarks emphasise single-goal theorem proving within static contexts, diverging significantly from the demands of real-world proof maintenance.
+As shown in Table 1, conventional benchmarks evaluate success based on isolated proof completion, neglecting the broader engineering workflows needed to evolve and maintain large formal mathematics libraries.
+Recent efforts have explored richer proof generation scenarios [4, 6, 11, 12, 5, 13, 1], but still typically focus on isolated reasoning tasks rather than realistic proof engineering workflows.
+A clear gap remains in systematic, task-diverse benchmarks that reflect the full scope of real-world proof engineering.
 Aspect Current Benchmarks Real-world Practice
 Scope Isolated, static theorem proving Dynamic, evolving large libraries
-Objectives Single-attempt theorem completion Iterative refinement, multiple engineering goals
-Activities Individual proof completion Feature addition, proof refactoring, bug fixing
-Table 1: Comparison between conventional theorem-proving benchmarks and real-world proof engineering practices.
-Several recent efforts have explored richer proof generation scenarios, such as incorporating intermediate steps, leveraging auxiliary definitions [6, 9, 15, 16, 7], or modifying existing proofs through refactoring [18] and improvement techniques [1].
-However, these approaches are still typically used as tools for data augmentation or pipeline enhancement, rather than serving as rigorous evaluation frameworks.
-As a result, a clear gap remains in the form of systematic, task-diverse benchmarks that reflect the full scope of real-world proof engineering.
+Objectives Individual proof completion Feature addition, proof refactoring, bug fixing
+Table 1: Comparison between conventional benchmarks and real-world proof engineering practices.
 3 APE-Bench I: A Realistic Benchmark for File-Level Proof Engineering
-To concretely realize the paradigm of Automated Proof Engineering (APE), we introduce APE-Bench I, a realistic and instruction-driven benchmark constructed from real-world edits in the Mathlib4 formal mathematics library.
-This section describes the design and construction of the benchmark, including its task format, the methodology for extracting and filtering file-level edits from version history, and the automatic synthesis of natural-language instructions and semantic labels.
-Together, these components enable structured, modular, and semantically faithful evaluation of language models on localized proof engineering tasks.
+To concretely realise the paradigm of Automated Proof Engineering (APE), we introduce APE-Bench I, a realistic and instruction-driven benchmark constructed from real-world edits in the Mathlib4 formal mathematics library.
+This section describes the design and construction of the benchmark, including its task format, the methodology for extracting file-level edits from version history, and the automatic synthesis of natural-language instructions and semantic labels.
 3.1 Task Format and Objective
-Each task in APE-Bench I defines a localized file-level proof engineering problem, grounded in an actual edit to a Lean source file.
-The objective is to apply a valid patch that fulfills a natural-language instruction by transforming a given pre-edit file into a structurally correct and semantically meaningful post-edit version.
+Each task in APE-Bench I defines a localised file-level proof engineering problem, grounded in an actual edit to a Lean source file.
+The objective is to apply a valid patch that fulfils a natural-language instruction by transforming a given pre-edit file into a structurally correct and semantically meaningful post-edit version.
 […]
@@
   • PreFile contains the complete Lean source code of the target file before the edit. It provides the context necessary to interpret the instruction and generate a corresponding structural change.
-  • Patch is a unified diff that encodes the intended edit, specifying insertions and deletions relative to PreFile.
-The diff format supports modular and interpretable editing, and facilitates automated application, verification, and evaluation.
-This triplet formulation reflects the real-world dynamics of proof maintenance, where developers operate on full files with evolving structures rather than isolated subgoals.
-It also aligns naturally with instruction-following language models, enabling instruction-conditioned generation and evaluation of file-level structural changes.
-By grounding each task in a concrete diff from version history, APE-Bench I preserves the realism, coherence, and modularity essential for evaluating automated proof engineering capabilities.
-3.2 Task Extraction from Mathlib Commits
-(a) Token distribution by benchmark
+  • GoldDiff is a unified diff extracted from human-written commits that serves as the task’s ground truth, specifying changes relative to PreFile.
+Its format enables modular editing and supports automated application, verification, and evaluation.
+This triplet formulation mirrors the practical realities of proof maintenance, wherein developers continuously update and refactor multiple interdependent components within full files, as opposed to resolving isolated subgoals in static settings.
+3.2 Task Extraction from Mathlib4 Commits
+(a) Token distribution by benchmarks
 (b) Candidate reduction across filtering stages
 Figure 3: APE-Bench I: Scale and complexity of extracted tasks.
-(a) Compared to prior datasets, APE-Bench I tasks span significantly longer input–output contexts, reflecting greater structural and semantic complexity.
-(b) Starting from 19k commits, the extraction pipeline produces over 185k candidates, which are refined through multiple filtering stages to yield 10,928 high-quality tasks.
-To construct realistic and structurally diverse proof engineering tasks, we extract file-level edits from the version history of Mathlib4, Lean’s primary formal mathematics library.
-The resulting benchmark tasks are derived from actual developer activity, ensuring contextual fidelity and development relevance.
-This section describes the extraction pipeline, which proceeds in four stages: commit sourcing and decomposition, chunk-level expansion, multi-stage filtering, and syntactic validation.
+(a) APE-Bench I tasks span significantly longer input–output contexts than prior datasets.
+(b) Starting from 19k commits, the pipeline produces 185k+ candidates, filtered to 10,928 high-quality tasks.
+To construct realistic and diverse proof engineering tasks, we extract file-level edits from the version history of Mathlib4.
+Our extraction pipeline, illustrated in Figure 3(b), proceeds in four stages:
 Step 1: Commit Sourcing and Diff Decomposition.
-We begin with 19,867 Git commits made between August 2023 and March 2025, a period selected to capture stable post-migration development in Mathlib4.
+We begin with 19,867 Git commits made between August 2023 and March 2025.
 Each commit is decomposed into file-level diffs, isolating modified files and their corresponding changes.
-These diffs are further segmented into semantically scoped chunks: for modified files, we use hunk-level boundaries; for newly added files, we apply heuristics based on Lean declaration structure to determine logical segments.
+These diffs are further segmented into semantically scoped chunks by hunk or declaration boundaries.
 Step 2: Granular Chunk Expansion.
-To improve modeling tractability and enhance semantic interpretability, we decompose large diffs into smaller, contiguous edit fragments.
-Each fragment is applied incrementally to the evolving file state, yielding fine-grained and contextually grounded task units.
-This expansion increases the total number of edit candidates by nearly an order of magnitude, producing over 187,000 candidate units from the original commit set.
+To improve modelling tractability and enhance semantic interpretability, we decompose large diffs into smaller, contiguous edit fragments.
+Each fragment is applied incrementally to the evolving file state, yielding fine-grained task units.
+This expansion increases the total number of edit candidates to over 187,000.
 Step 3: Multi-Stage Filtering.
-To isolate high-quality, semantically meaningful tasks, we apply a sequence of automated filters targeting three dimensions:
-  • Semantic relevance: We retain only commits with recognizable editing intent, identified via conventional prefixes such as feat, fix, or refactor.
-Commits involving file deletions, mechanical formatting, or edits outside the Mathlib/ directory are excluded.
-  • Content quality: Edits that only affect comments, import ordering, or whitespace are discarded, as they lack structural substance.
-  • Structural integrity: We restrict each task to at most 100 lines of non-comment changes and a maximum input length of 16,000 tokens (measured using the GPT-4o tokenizer).
-Tasks must be localized, coherent, and semantically self-contained.
-The most substantial reduction occurs after chunking, where structurally trivial diffs are eliminated.
-Subsequent filters further refine the corpus.
-The overall retention trajectory is visualized in Figure 3(b), illustrating each pruning stage in the pipeline.
+We apply a sequence of automated filters targeting semantic relevance (retaining only commits with recognisable editing intent), content quality (discarding comment-only or whitespace edits), and structural integrity (restricting to at most 100 lines of changes and 16,000 tokens).
+The overall retention trajectory is visualised in Figure 3(b).
+For additional details on our filtering methodology, see Appendix B.1.
 Step 4: Compilation Validation.
-To ensure structural soundness, all retained patches are compiled under their post-edit environments using Lean.
-Only those patches that successfully typecheck in the restored version context are included in the final benchmark.
-This step guarantees that each task is grounded in a valid and verifiable edit to a Lean file.
-The final benchmark consists of 10,928 structurally coherent file-level tasks.
-As shown in Figure 3(a), these tasks exhibit significantly longer input and output sequences than prior benchmarks such as miniF2F and miniCTX.
-This reflects the higher structural complexity of proof engineering tasks and highlights the need for long-context reasoning and precise structural planning.
-Each stage of the pipeline incrementally refines the dataset to ensure structural integrity, semantic depth, and fidelity to real-world development practices.
-The resulting benchmark comprises 10,928 structurally coherent file-level tasks that capture a diverse set of realistic, non-trivial modifications authored by human contributors.
-This dataset serves as the foundation for instruction synthesis and stratified evaluation in the subsequent sections.
-As shown in Figure 3(a), these tasks exhibit significantly longer input and output sequences than prior benchmarks such as miniF2F and miniCTX, underscoring the greater structural complexity of proof engineering and the need for long-context reasoning and precise structural planning.
+To ensure syntactic soundness, all retained patches are compiled under their post-edit environments using Lean.
+Only patches that are successfully compiled in the restored version context are included in the final benchmark.
+The final benchmark consists of 10,928 structurally coherent tasks.
+As shown in Figure 3(a), these tasks exhibit significantly longer input and output sequences than prior benchmarks, reflecting the higher structural complexity of proof engineering tasks.
 3.3 Instruction Synthesis and Task Labelling
-To enable instruction-driven modeling and facilitate stratified evaluation, each task in APE-Bench I is paired with a natural language instruction and annotated with semantic labels describing its functional category and structural difficulty.
-Both instruction synthesis and task labelling are performed automatically using LLMs, with verification procedures designed to ensure fidelity, clarity, and modeling relevance.
-Instruction Generation.
-For each filtered task derived from a Lean file diff (as described in Section 3.2), we synthesize an accompanying instruction that concisely captures the developer’s intent.
-We use Claude Sonnet 3.7 (thinking mode) to generate instructions from the raw diff, prompting the model to extract and summarize the intended edit in an imperative, concise, and actionable style.
-Instructions typically reflect changes such as introducing a new definition, modifying a proof, or restructuring an existing theorem.
-To ensure semantic alignment between instruction and patch, we apply a reverse validation procedure identical to the semantic evaluation process described in Section 4.2.
-Specifically, the gold patch is re-evaluated against the synthesized instruction using the LLM-as-a-Judge framework.
-Only tasks for which the original edit satisfies the instruction are retained in the benchmark.
-This backward verification ensures that instructions are both meaningful and faithfully representative of the structural change.
-Task Labelling.
-Figure 4: Distribution of code change magnitudes across semantic labels within the test set.
-(a) Based on non-comment line counts from gold patches in the test set, grouped by difficulty.
-(b) Patch sizes grouped by functional category, based on gold commit diffs.
-Figure 5: Object-level statistics of Lean patches in the test set. Distribution of added and changed def and theorem declarations, measured by count and total lines.
-To enable stratified evaluation and reveal model capability differences across structurally diverse tasks, each instance in the test set is automatically annotated with semantic labels describing its functional intent and structural complexity.
-Specifically, we assign three types of labels:
-  • Task Category, capturing functional purpose, is categorized as Feature, Refactor, or Bug Fix, aligned with common proof engineering operations.
-  • Difficulty Level, estimating structural and reasoning complexity, is stratified as Very Easy, Easy, Medium, Hard, or Very Hard.
-  • Task Nature, a binary label indicating whether a patch is Substantial or Superficial, used for filtering; only Substantial tasks are retained in the benchmark.
-In addition to supporting dataset annotation, the labels also guide task selection: instances marked as Very Easy or Superficial are filtered out, ensuring that the benchmark retains only structurally substantive and semantically meaningful edits.
-Beyond this filtering role, the labels enable stratified evaluation across both functional categories and complexity levels, forming the foundation for the fine-grained performance analysis in Section 5.3.
-We next assess the effectiveness of these labels by analyzing their correlation with code magnitude and declaration-level structure.
-Statistical Validation of Difficulty Labels.
-Figure 4(a) shows that while Easy tasks are generally small, Medium and Hard tasks span a broad range of edit sizes, with no strict separation based on line count.
-This indicates that difficulty labels capture more than superficial metrics—they reflect underlying semantic and structural complexity.
-This is further confirmed by object-level patterns in Figure 5.
-Medium and Hard tasks involve more declarations and larger edits per object, particularly for theorems, suggesting greater logical coordination and proof effort.
-These trends align with empirical results in Section 5.3, where model performance declines sharply from Easy to Hard, validating the semantic significance of the difficulty labels.
-Functional Differentiation by Task Category.
-Task categories also reveal distinct structural roles.
-As shown in Figure 4(b), Feature tasks are larger and often involve new declarations, while Refactor and Bug Fix edits are smaller and focus on modifying existing code.
-Object-level statistics support this: Feature tasks add theorems and definitions, while the other two primarily adjust existing defs.
-These differences reflect practical development workflows.
-The small scale of Refactor and Bug Fix edits stems in part from filtering heuristics, which exclude trivial or mechanical changes.
-As such, retained tasks across all categories represent meaningful engineering actions.
+Figure 4: Distribution of code change magnitudes across semantic labels in the test set.
+(a) Based on non-comment line counts from gold patches, grouped by difficulty.
+(b) Patch sizes grouped by functional category.
+Each task is paired with a natural language instruction and annotated with semantic labels describing its functional category and structural difficulty.
+For instruction generation, we use Claude Sonnet 3.7 (thinking mode) to extract and summarise the developer’s intent from the raw diff in an imperative style.
+We apply a reverse validation procedure to ensure semantic alignment between instruction and patch (detailed methodology in Appendix B.2).
+For task labelling, we assign three types of labels to the test set:
+  • Task Category: Feature, Refactor, or Bug Fix, aligned with common proof engineering operations.
+  • Difficulty Level: Very Easy, Easy, Medium, Hard, or Very Hard, reflecting increasing structural and logical complexity. Very Easy tasks are filtered out.
+  • Task Nature: Substantial or Superficial, used for filtering; only Substantial tasks are retained.
+Figure 4 shows the distribution of code changes across difficulty levels and functional categories.
+These labels enable stratified evaluation and help reveal capability gaps across models, as confirmed by our experimental results in Section 5.
+Appendix B.3 provides additional statistical validation of these label distributions.
 4 Evaluation Infrastructure and Protocol
+To enable reliable assessment of model-generated edits in APE-Bench I, we establish a two-stage evaluation framework that reflects the dual demands of proof engineering: syntactic correctness and semantic adequacy.
+As illustrated in Figure 5, each candidate patch is evaluated for both grammatical validity under Lean’s type system and alignment with the task’s instruction-level intent.
+The first stage, syntactic verification, checks whether the resulting file passes Lean verification within the corresponding versioned environment.
+This process is implemented via Eleanstic, a scalable verification infrastructure that supports reproducible, parallelisable, and version-aware Lean verification (Section 4.1).
+The second stage, semantic judgement, determines whether the patch satisfies the instruction as intended.
+This is carried out by a large language model acting as an instruction-conditioned judge (Section 4.2).
+A task is marked as successful only if a model-generated patch passes both verification stages.
+This protocol provides a standardised and scalable basis for computing success metrics, enabling detailed analyses across syntax-level and semantics-level dimensions.
+4.1 Syntactic Verification via Eleanstic
 Figure 5: Evaluation pipeline for APE-Bench I.
-A model-generated patch undergoes syntactic verification via the Lean compiler and semantic judgement via an LLM-as-a-Judge.
-A task is marked as successful only if both stages are passed.
-To enable reliable and automated assessment of model-generated edits in APE-Bench I, we establish a two-stage evaluation framework that reflects the dual demands of proof engineering: syntactic correctness and semantic adequacy.
-Each candidate patch is evaluated for both grammatical validity under Lean’s type system and alignment with the task’s instruction-level intent.
-The first stage, syntactic verification, checks whether the resulting file of the generated pass the Lean verification within the corresponding versioned environment.
-This process is implemented via Eleanstic, a scalable verification infrastructure that supports reproducible, parallelizable, and version-aware Lean verification infrastructure (Section 4.1).
-The second stage, semantic judgement, determines whether the patch satisfies the instruction as intended.
-This is carried out by a large language model acting as an instruction-conditioned judge, which reasons over the pre-edit file, the patch, and the goal description to determine task success (Section 4.2).
-Figure 6 presents the overall evaluation pipeline.
-A task is marked as successful only if a model-generated patch passes both verification stages.
-This protocol provides a standardized, interpretable, and scalable basis for computing success metrics, and enables detailed analysis of model performance across syntax-level and semantics-level axes.
-4.1 Syntactic Verification via Eleanstic
+Given a commit hash, the corresponding Lean environment is restored by a snapshot from content-addressable storage (CAS), reproducing the exact worktree state.
+An LLM generates a patch based on the instruction and the original file.
+This patch is applied to a temporary file and evaluated through Lean compile and LLM judge.
 Syntactic verification in APE-Bench I requires compiling model-generated patches within the exact versioned environment from which each task was derived.
-In large-scale formal libraries like Mathlib4, where definitions and dependencies evolve rapidly across commits, such version fidelity is critical to ensure meaningful evaluation.
-However, this fidelity introduces significant engineering complexity: naively rebuilding the library for every historical commit results in both prohibitive computational cost and unsustainable storage overhead.
-A single clean build of Mathlib4 can take several hours to complete and produce multiple gigabytes of intermediate artifacts.
-When scaled to thousands of commits—each potentially associated with thousands of patch candidates and evaluated across multiple models—this cost compounds rapidly, impeding both throughput and reproducibility.
-To address this, we introduce Eleanstic, a dedicated infrastructure for scalable, version-aware syntactic verification.
-Eleanstic decouples the high-cost, stateful process of building Mathlib from the low-cost, stateless process of validating edits.
-It does so by collapsing each post-build state into a compact, content-addressable snapshot that can later be restored to yield a faithful Lean environment for verification.
-This architecture significantly reduces redundant computation, minimizes on-disk storage usage, and supports efficient parallel evaluation across thousands of distinct versions.
-Snapshot Generation.
-For each unique commit, Eleanstic first performs a clean lake build and recursively traverses the resulting worktree.
-Every file is hashed using SHA-256 and stored in a content-addressable storage (CAS) structure with a two-level hash-based directory layout.
-In parallel, a binary snapshot is created to record the mapping from relative file paths to content hashes and file types (regular or symbolic).
-This snapshot, encoded using fixed-width binary records, serves as a complete and portable blueprint of the post-build environment, ensuring reproducibility while minimizing metadata overhead.
-Snapshot Restoration.
-At verification time, the appropriate snapshot is loaded and its contents restored into a temporary, isolated worktree that exactly replicates the post-build file structure of the original compilation.
-Files are fetched from the global CAS and written to disk.
-This process is stateless and fully parallelizable, enabling fast switching across historical commits without requiring repeated builds.
-Empirically, Eleanstic restores a full Lean environment in under one second on SSD-backed systems.
-To support long-term scalability, Eleanstic performs aggressive global deduplication across all snapshots.
-Across 3734 commits spanning August 2023 to March 2025, it compresses 15.6 TB of build artifacts into just 1.1 TB, yielding a storage reduction of over 93%.
-This compression enables persistent caching of every historical state required for evaluation, without exhausting disk resources.
-Patch Verification.
-After restoring the environment, the model-generated patch (see Section 5.1) is applied to the target Lean file, and the resulting content is saved to a temporary file.
-Verification follows a strict file-level isolation policy: rooted in the restored Mathlib worktree, the Lean compiler compiles only the temporary file without affecting any files in the restored environment.
-Additionally, any temporary file that attempts to import the target file is discarded to prevent unintended leakage across file boundaries.
-A patch is considered syntactically valid if it compiles successfully within the corresponding version of Mathlib4 and does not emit warnings.
+In large-scale formal libraries like Mathlib4, where definitions and dependencies evolve rapidly across commits, such version fidelity is critical for meaningful evaluation.
+We introduce Eleanstic, a dedicated infrastructure for scalable, version-aware syntactic verification.
+Eleanstic decouples the high-cost process of building Mathlib4 from the low-cost process of validating edits by collapsing each post-build state into a compact, content-addressable snapshot.
+This architecture significantly reduces redundant computation, minimises storage usage, and supports efficient parallel evaluation across thousands of distinct versions (technical implementation details in Appendix B.4).
+For verification, a model-generated patch is applied to the target Lean file in an isolated environment, and the Lean compiler verifies only the modified file without affecting the restored environment.
+A patch is considered syntactically valid if it compiles successfully and does not emit warnings.
 4.2 Semantic Judgement via LLM-as-a-Judge
@@
 5.1 Experimental Setup
-We evaluate seven general-purpose language models on APE-Bench I using a standardized evaluation pipeline.
-The task for each model is to generate a Lean patch that fulfills a natural language instruction, based on a complete pre-edit source file.
-Success is measured via a two-stage evaluation combining syntactic verification and semantic judgement.
-Models.
-We evaluate a range of state-of-the-art, general-purpose LLMs, selected to represent both long-context reasoning and standard generation capabilities:
-  • Reasoning models: OpenAI o3-mini, Claude Sonnet 3.7 (thinking), DeepSeek R1.
-  • Non-reasoning models: Claude Sonnet 3.7, DeepSeek V3 (250324), GPT-4o, Doubao 1.5 Pro.
-All models are zero-shot, without any task-specific fine-tuning for proof engineering.
-Patch Normalization via DiffRepair.
-Model outputs are produced in unified diff format, which often contains structurally imprecise or misaligned edits (e.g., incorrect line anchors, out-of-date contexts).
-To ensure reliable application and verification, we introduce a post-processing phase called DiffRepair, which rewrites imprecise diffs into structurally consistent patches.
-This mechanism operates in two phases: first, it localizes the intended edit position using fuzzy anchor matching; then it reconstructs the patch with preserved structural and contextual integrity.
-All generated patches are normalized through DiffRepair before subsequent compilation or evaluation.
-(See Appendix A for full technical details.)
-Sampling Methodology.
-We adopt the standard pass@$k$ evaluation [4] to estimate task success rates. For each task, $n=20$ candidate patches are sampled, and pass@16 is computed using the unbiased estimator:
+We evaluate eight state-of-the-art language models on APE-Bench I: Gemini 2.5 Pro Preview, o3-mini, Claude Sonnet 3.7 (with and without thinking mode), DeepSeek R1, DeepSeek V3, GPT-4o, and Doubao 1.5 Pro.
+All models generate patches in unified diff format without task-specific fine-tuning.
+We process outputs using DiffRepair, a post-processing mechanism that rewrites imprecise diffs into structurally consistent patches (described in Appendix C.3).
+We adopt the standard pass@$k$ evaluation [3] to estimate task success rates.
+For each task, $n=20$ candidate patches are sampled, and pass@16 is computed using the unbiased estimator:
 […]
-where $c\leq n$ is the number of patches passing both syntactic and semantic verification.
-Temperature Settings.
-For models supporting temperature control, we evaluate outputs under two sampling temperatures, $T\in\{0.0,0.6\}$.
-Empirical results indicate that $T=0.6$ consistently improves output diversity and task coverage.
-For reasoning models that do not expose temperature parameters, we adopt default or recommended configurations: o3-mini is run with medium reasoning effort, while Claude Sonnet (thinking) is allocated 6,000 tokens of thinking budget to enable extended context comprehension.
+where $c\leq n$ is the number of patches passing both syntactic and semantic judgement. For most models, we use temperature T=0.6 to improve output diversity.
 5.2 Main Results
 We evaluate model performance using the pass@16 metric, which estimates the probability that at least one of 16 generated candidates satisfies specified criteria. Specifically, we distinguish between:
-  • Lean Compile pass@16: The proportion of tasks for which at least one candidate compiles successfully using Lean 4, validating syntactic and structural soundness.
-  • LLM-Judge pass@16: The proportion of tasks for which at least one compiled candidate additionally satisfies the instruction according to semantic evaluation via LLM-as-a-Judge.
-  • Semantic Drop: The percentage decrease from Lean Compile to LLM-Judge success, reflecting how often syntactically valid outputs fail to meet the intended edits.
-Table 2 summarizes the performance of seven general-purpose models.
-Among all candidates, o3-mini achieves the highest overall success rate, with 33.58% of tasks completed successfully under full evaluation.
-It also shows the highest syntactic pass rate at 48.38%, indicating its ability to generate compilable edits across a broad range of task types.
-However, its semantic drop-the proportion of compilable patches that fail to meet instruction-level requirements-is also substantial at over 30%, suggesting frequent deviation from intended modifications.
-In contrast, Claude Sonnet 3.7 (thinking) exhibits perfect semantic consistency with a semantic drop of zero, yet achieves a significantly lower overall success rate due to reduced syntactic coverage.
-GPT-4o and Doubao 1.5 Pro show the opposite trend: both suffer from large semantic drops of 49.31% and 32.21%, respectively.
-These patterns highlight fundamental differences in model behavior—while some models excel at generating well-formed Lean code, they often fail to capture the semantic intent of the editing instruction.
-Model Lean Compile pass@16 LLM-Judge pass@16 Semantic Drop (%)
-o3-mini 48.38 33.58 30.58
-Claude Sonnet 3.7 (thinking) 17.73 17.73 0.00
-DeepSeek R1 17.67 15.85 10.30
-DeepSeek V3 16.23 15.72 3.11
-Claude Sonnet 3.7 15.85 14.85 6.31
-GPT-4o 18.94 9.60 49.31
-Doubao 1.5 Pro 9.70 6.58 32.21
-Table 2: Overall model performance on APE-Bench I using pass@16. Semantic Drop is computed as the percentage of Lean-compilable outputs that fail LLM-based semantic validation.
-Figure 7 further illustrates model behavior across different pass@$k$ thresholds.
+  • Verification pass@16: The proportion of tasks for which at least one candidate compiles successfully using Lean 4.
+  • Judgement pass@16: The proportion of tasks for which at least one compiled candidate additionally satisfies the instruction according to semantic evaluation.
+  • Relative Decrease: The percentage decrease from Verification to Judgement success, reflecting how often syntactically valid outputs fail to meet the intended edits.
+Model Verification pass@16 Judgement pass@16 Relative Decrease (%)
+Gemini 2.5 Pro Preview 18.79% 18.04% 3.99%
+Claude Sonnet 3.7 (thinking) 12.73% 12.73% 0.001%
+DeepSeek R1 15.30% 12.55% 17.97%
+DeepSeek V3 12.44% 11.81% 5.06%
+Claude Sonnet 3.7 11.33% 10.83% 4.41%
+o3-mini 20.13% 8.60% 57.28%
+GPT-4o 15.16% 6.73% 55.61%
+Doubao 1.5 Pro 6.19% 2.98% 51.86%
+Table 2: Overall model performance on APE-Bench I using pass@16. Relative Decrease is computed as the percentage of Lean-compilable outputs that fail LLM-based semantic judgement.
+As shown in Table 2, Gemini 2.5 Pro Preview achieves the highest judgement success rate at 18.04%, while o3-mini leads in verification pass rate (20.13%) but suffers from a substantial relative decrease (57.28%).
+Claude Sonnet 3.7 (thinking) exhibits nearly perfect semantic consistency with a negligible relative decrease (0.001%), though at a lower overall success rate.
+Notably, even the best-performing model succeeds on only about one-sixth of the benchmark tasks, highlighting the substantial gap between present capabilities and the demands of real-world proof engineering.
+Additional performance metrics across different pass@k values are available in C.
+5.3 Structural Robustness and Task-Scale Sensitivity
+Figure 6: Model pass@16 success rate grouped by ground-truth patch size.
+Performance declines rapidly with increasing patch size.
+Gemini 2.5 Pro Preview leads in small to medium ranges, but no model demonstrates robustness beyond 50-line edits.
+To assess the structural robustness of current models, we group benchmark tasks by the size of their ground-truth patches and evaluate pass@16 success rates within each bin.
+As shown in Figure 6, most models experience a sharp drop in performance as the number of modified lines increases.
+On small-scale edits (1–5 lines), several models achieve moderate to high success rates.
+However, performance deteriorates significantly on medium (11–20 lines) and large-scale edits (>20 lines), with all models collapsing to near-zero success beyond 50 lines.
+Gemini 2.5 Pro Preview consistently performs best in the low to medium complexity range, but fails to generalize to structurally demanding scenarios.
+These results reveal a fundamental scalability challenge: current LLMs are not robust to increasing structural complexity in proof engineering tasks.
+Figure 7: Heatmap of evaluation pass@16 success rates across task types, difficulty levels, and mathematical modules.
+Column widths represent the proportion of each class in the test set.
+Color intensity indicates model performance.
+We further analyze model performance across task category, difficulty level, and mathematical domain. Figure 7 presents a unified heatmap summarizing pass@16 success rates across these axes.
+Task Category.
+Models generally handle Bug Fix tasks better than other categories, likely due to their localized nature and relatively narrow scope.
+Refactor tasks reveal greater variation: Gemini and Claude variants achieve relatively strong performance, benefiting from architectural biases toward structural preservation, while GPT-4o and Doubao consistently underperform.
+Feature addition is the most challenging category across all models, requiring the introduction of new formal elements while maintaining global coherence.
+Difficulty Level.
+A steep capability cliff is evident across difficulty levels.
+Most models perform adequately on Easy tasks, but success rates fall sharply for Medium tasks and approach zero on Hard tasks.
+These harder tasks typically involve extensive edits, multi-step dependencies, or cross-declaration modifications, indicating a fundamental weakness in current models’ ability to reason over large or interdependent proof contexts.
+Mathematical Domain.
+Domain-level performance is also highly uneven.
+Tasks from Dynamics and Tactic modules are consistently easier for all models, likely due to simpler proof structures and shorter edit spans.
+In contrast, tasks from Combinatorics, Measure Theory, Probability, and Set Theory show uniformly low success rates, reflecting their greater formal density and structural complexity.
+Taken together, these results indicate that current LLMs lack structural robustness and task-scale adaptability.
+While they perform well on localized and syntactically constrained tasks, their effectiveness drops sharply on edits requiring broader semantic planning and structural integration.
+Overcoming this limitation will require models to incorporate stronger edit planning, context understanding, and global coordination mechanisms, potentially through agentic or feedback-driven architectures.
+5.4 Semantic Robustness and Failure Characterization
+Error Class Description
+Component Omission Required definitions, proofs, or examples are missing
+Incompleteness Issue The edit is structurally partial, unfinished, or placeholder
+Mathematical Error The logic is invalid or inconsistent with Lean’s formal semantics
+Quality Defect Redundant or unabstracted edits, or lack of documentation
+Requirement Mismatch Misalignment with the intended instruction
+Table 3: Semantic error classes and their typical manifestations.
+Figure 8: Distribution of semantic error types across models (log-scaled y-axis).
+Component omission, quality defects, and incompleteness are the most frequent error types.
+Claude Sonnet 3.7 (thinking) demonstrates consistently minimal error rates across all categories.
+To investigate the gap between syntactic validity and semantic correctness, we conduct a systematic analysis of failure cases using a taxonomy of five high-level semantic error classes, summarized in Table 3.
+These categories capture recurring patterns where models produce syntactically valid but semantically inadequate patches.
+Error Distribution.
+Figure 8 illustrates the frequency distribution of semantic error types across different models.
+Component omission emerges as the most common failure mode, where models neglect to include essential definitions, theorems, or examples required to fulfill the task.
+This is followed by quality defects, such as redundant or unrefined edits and missing documentation, and incompleteness issues, where edits are only partially implemented or left unfinished.
+Model Differences.
+GPT-4o and o3-mini exhibit particularly high error rates across all categories, often producing outputs that are not only mathematically incorrect but also misaligned with the original instruction.
+These models frequently fail to plan edits holistically, leading to missing logic or irrelevant changes.
+In contrast, Claude Sonnet 3.7 (thinking) shows the lowest error frequency in every category, reflecting strong semantic robustness, structural alignment, and adherence to instructions.
+The analysis reveals that semantic failures are not idiosyncratic but follow consistent structural patterns.
+Most stem from limited reasoning depth, lack of multi-step planning, or failure to reconcile edits with the broader file context.
+These deficiencies persist even when syntactic correctness is achieved, underscoring the insufficiency of type-checking alone as a proxy for task completion.
+5.5 Judge Model Selection and Evaluation Bias
+Model Evaluated
+Claude
+Thinking (%)
+Gemini (%)
+DeepSeek
+R1 (%)
+DeepSeek
+V3 (%)
+Majority
+Vote (%)
+Gemini 2.5 Pro Preview 18.04 +0.55 16.20 -1.29 17.66 +0.17 18.05 +0.56 17.49
+DeepSeek R1 12.55 -0.38 10.56 -2.37 13.80 +0.87 14.80 +1.87 12.93
+Claude Sonnet 3.7 (thinking) 12.73 +0.31 11.47 -0.95 12.73 +0.31 12.73 +0.31 12.42
+o3-mini 8.60 -2.22 5.94 -4.88 15.31 +4.49 13.42 +2.60 10.82
+Claude Sonnet 3.7 10.83 +0.25 8.84 -1.74 11.33 +0.75 11.33 +0.75 10.58
+DeepSeek V3 11.81 +0.53 9.44 -1.84 11.93 +0.65 11.92 +0.64 11.28
+GPT-4o 6.73 -1.85 5.71 -2.87 13.31 +4.73 8.58 0.00 8.58
+Doubao 1.5 Pro 2.98 -0.68 3.48 -0.18 4.08 +0.42 4.08 +0.42 3.66
+Squared Deviation from Majority 3.31 78.02 233.61 81.69 -
+Family Preference Index (FPI) +0.95% +2.12% –1.16% +1.10% -
+Table 4: Judge model comparison: Success rates across different LLM judges.
+Each judge column shows score and deviation from majority (colored: green for positive, red for negative).
+The bottom rows report total deviation from consensus (squared) and the Family Preference Index (FPI), which quantifies each judge’s average relative bias toward its own model family.
+While syntactic validity can be objectively determined by the Lean compiler, semantic evaluation requires subjective judgment.
+In our framework, this is delegated to large language models acting as instruction-conditioned judges.
+However, since these judges are themselves LLMs, a natural concern arises: do they exhibit systematic bias, particularly favoring outputs from models within their own family?
+To measure this effect, we introduce a Family Preference Index (FPI), which quantifies the relative deviation between a judge’s ratings of same-family versus other-family models.
+Formally, for judge $J$,
+$\text{FPI}_{J}=\mathbb{E}_{m\in M_{\text{same}}}[\Delta_{J,m}]-\mathbb{E}_{m\in M_{\text{other}}}[\Delta_{J,m}]$
+where $\Delta_{J,m}$ is the deviation between $J$’s score and the majority consensus for model $m$. A positive FPI suggests preferential scoring toward same-family outputs.
+As summarized in Table 4, most judges display mild but measurable family preferences.
+Gemini and DeepSeek V3 exhibit positive FPIs, indicating upward bias for their own models.
+Claude Sonnet 3.7 (thinking) shows the smallest FPI and the lowest deviation from majority voting, reflecting high consistency and relative neutrality.
+In contrast, DeepSeek R1 reveals a reversed pattern, assigning more favorable ratings to non-family outputs.
+While no judge displays extreme favoritism—most FPI values remain within ±2%—such biases can subtly influence evaluations, especially when models are closely matched.
+These findings motivate our choice of Claude Sonnet 3.7 (thinking) as the primary semantic evaluator, based on its empirical alignment with consensus and minimal systemic bias.
+To further improve robustness, future work may incorporate multi-judge aggregation, adversarial calibration, or dedicated reference evaluators.
+6 Conclusion
+This work introduces the paradigm of Automated Proof Engineering (APE) and presents APE-Bench I, the first large-scale benchmark grounded in real-world proof development workflows from Mathlib4.
+By reframing formal mathematics as instruction-driven code editing and combining Lean compilation with LLM-based semantic evaluation, APE-Bench I offers a rigorous framework for evaluating LLMs in realistic proof engineering tasks.
+Our empirical results reveal significant performance gaps among state-of-the-art models, particularly under structurally complex scenarios.
+A key limitation of APE-Bench I is its focus on single-file edits.
+Future work will extend this to APE-Bench II and III, which will address multi-file coordination, project-scale verification, and ultimately, autonomous agents capable of planning, editing, and repairing formal libraries.
+This benchmark series lays the foundation for developing next-generation systems for practical and scalable formal proof maintenance.
+Appendix A Extended Background and Related Work
+A.1 Formal Proof Engineering: Practice and Challenges
+Real-world formal verification projects, including Lean’s Mathlib [9], Isabelle’s Archive of Formal Proofs, and the seL4 microkernel verification project [8], exemplify the scale and complexity of proof engineering.
+These libraries often contain thousands of interdependent definitions, theorems, and proofs, where modifying a single element may trigger cascading changes across the codebase.
+Ensuring logical consistency under such conditions demands significant manual labour and deep structural awareness.
+While certain tools-such as iCoq for regression-proving [2]-provide partial automation, they are largely heuristic and require intensive human oversight.
+To date, there remains a lack of automated systems capable of performing high-fidelity edits in complex proof environments.
+A.2 Limitations of Existing Benchmarks
+The current gap between theorem-proving benchmarks and practical proof engineering extends beyond the aspects highlighted in Table 1.
+Most benchmarks focus on mathematical reasoning skills in isolation, but lack the context of evolving codebases, dependency management, and continuous integration requirements common in real-world libraries.
+Several recent efforts have explored richer proof generation scenarios, such as incorporating intermediate steps, leveraging auxiliary definitions [4, 6, 11, 12, 5], or modifying existing proofs through refactoring [13] and improvement techniques [1].
+However, these approaches are still typically used as tools for data augmentation or pipeline enhancement, rather than serving as rigorous evaluation frameworks that capture realistic development workflows.
+A.3 Comparison with Software Engineering Benchmarks: The Case of SWE-Bench
+The evolution towards more realistic benchmarks is a shared trend across automated reasoning domains.
+In software engineering, the introduction of benchmarks like SWE-Bench [7] marked a significant step, moving from isolated code generation tasks to evaluating LLMs on resolving real-world GitHub issues within established Python repositories.
+SWE-Bench’s reliance on existing unit test suites for verifying functional correctness provides an objective and automated measure of success, offering a valuable paradigm.
+APE-Bench draws inspiration from this advancement, aiming to create a similarly realistic and challenging environment for automated proof engineering.
+Like SWE-Bench, it sources tasks from real-world commit histories and focuses on instruction-driven modifications.
+However, the distinct nature of formal mathematics compared to general software engineering necessitates adaptations in task structure and evaluation.
+Table 5 outlines the key distinctions.
+Aspect SWE-Bench (Software Eng.) APE-Bench (Proof Eng.)
+Core Task Resolve functional issues Implement proof modifications
+Syntactic Check Compiler/Interpreter Lean Compiler (type checking)
+Semantic Verification Unit tests for functional correctness LLM-as-a-judge for semantic alignment with instructions
+Objects Observable behavior, I/O correctness Logical validity, structural integrity, adherence to mathematical intent
+Table 5: Comparison between SWE-Bench and APE-Bench.
+A primary divergence lies in semantic evaluation.
+While software functionality can often be robustly verified by unit tests capturing desired behaviors, the semantic correctness of a proof modification is more nuanced.
+A syntactically valid proof (i.e., one that typechecks via the Lean compiler, analogous to code compilation) may still fail to accurately capture the intended mathematical argument, correctly implement a refactoring goal, or properly introduce a new definition as per the natural language instruction.
+The abstract logical structures and the specific intent behind a proof change are not always fully verifiable by pre-existing automated tests in the same direct manner.
+This distinction motivates APE-Bench’s hybrid evaluation strategy: combining compiler-based syntactic verification with an LLM-as-a-Judge (Section 4.2) to assess semantic fulfillment of the natural language instruction.
+This approach is a pragmatic response to the unique challenges of evaluating proof engineering tasks.
+Ultimately, APE-Bench aspires to evolve into a standard environment for developing and benchmarking increasingly sophisticated proof engineering agents, mirroring the role SWE-Bench plays in advancing agentic capabilities for software engineering.
+Appendix B Technical Details
+B.1 Detailed Commit Mining Methodology
+Our benchmark construction pipeline incorporates several specialised heuristics to ensure task quality and diversity:
+Semantic Filtering Implementation.
+To identify commits with recognisable intent, we analyse commit messages using regular expressions that match conventional prefixes like feat:, fix:, or refactor:.
+We also exclude commits that alter more than 10 files, as these typically represent mechanical changes rather than substantive engineering.
+Additionally, we filter out commits that only affect non-Mathlib/ directories (e.g., documentation or build scripts).
+Content Quality Metrics.
+Beyond simple line counting, we implement structural content analysis to identify semantically trivial changes:
+  • Whitespace ratio: We exclude diffs where >80% of changes affect only whitespace or empty lines.
+  • Comment-only changes: We detect and filter diffs that modify exclusively comment blocks (using Lean’s comment syntax patterns).
+  • Import statement changes: Changes affecting only import order or organisation are excluded.
+Chunk Segmentation Strategies.
+When segmenting modified files, we employ syntax-aware boundaries:
+  • For existing files, we use Lean’s declaration structure to identify logical boundaries.
+  • For newly added files, we use heuristics based on blank lines and declaration keywords.
+  • When expanding large diffs, we maintain declaration integrity by never splitting across def, theorem, or other key structural elements.
+These specialised techniques ensure that the resulting benchmark captures meaningful, self-contained, and structurally coherent proof engineering tasks rather than trivial or fragmented edits.
+B.2 Instruction Synthesis Methodology
+For instruction generation, we employ a carefully engineered prompt that directs Claude Sonnet 3.7 (thinking mode) to analyse a given diff and extract the underlying developer intent.
+The prompt specifically targets:
+  • Intent Extraction: Identifying the core purpose of the edit by analysing both structural changes and semantic modifications.
+  • Imperative Formulation: Reframing the extracted intent as a clear, actionable instruction using imperative voice.
+  • Technical Precision: Maintaining mathematical and Lean-specific terminology where appropriate.
+  • Scoping Accuracy: Ensuring the instruction’s scope corresponds to the actual changes in the diff.
+To validate instruction quality, we implement a reverse verification procedure where another LLM instance evaluates the alignment between the generated instruction and the actual diff.
+Instructions that do not sufficiently specify the target change or contain ambiguities are regenerated until they meet quality thresholds.
+B.3 Statistical Validation of Task Labels
+Figure 9: Object-level statistics of Lean patches in the test set. Distribution of added and changed def and theorem declarations, measured by count and total lines.
+Difficulty Labels Beyond Line Counts.
+Figure 4(a) shows that while Easy tasks are generally small, Medium and Hard tasks span a broad range of edit sizes, with no strict separation based on line count.
+This indicates that difficulty labels capture more than superficial metrics-they reflect underlying semantic and structural complexity.
+This is further confirmed by object-level patterns in Figure 9.
+Medium and Hard tasks involve more declarations and larger edits per object, particularly for theorems, suggesting greater logical coordination and proof effort.
+These trends align with empirical results in Section 5, where model performance declines sharply from Easy to Hard, validating the semantic significance of the difficulty labels.
+Functional Category Distinctions.
+Task categories also reveal distinct structural roles.
+As shown in Figure 4(b), Feature tasks are larger and often involve new declarations, while Refactor and Bug Fix edits are smaller and focus on modifying existing code.
+Object-level statistics support this: Feature tasks add theorems and definitions, while the other two primarily adjust existing defs.
+These differences reflect practical development workflows.
+The small scale of Refactor and Bug Fix edits stems in part from filtering heuristics, which exclude trivial or mechanical changes.
+As such, retained tasks across all categories represent meaningful engineering actions.
+B.4 Eleanstic: Version-Aware Verification Infrastructure
+Eleanstic provides a scalable, version-aware verification system for Lean code through a sophisticated snapshot mechanism. Its implementation consists of three key components:
+Snapshot Generation.
+For each unique commit, Eleanstic first performs a clean lake build and recursively traverses the resulting worktree.
+Every file is hashed using SHA-256 and stored in a content-addressable storage (CAS) structure with a two-level hash-based directory layout.
+In parallel, a binary snapshot is created to record the mapping from relative file paths to content hashes and file types (regular or symbolic).
+This snapshot, encoded using fixed-width binary records, serves as a complete and portable blueprint of the post-build environment, ensuring reproducibility while minimising metadata overhead.
+Our implementation runs on a single server with 128 CPU cores and 1TB of RAM.
+The initial compilation and snapshot generation for all 3734 commits from the Mathlib4 history required approximately 2 days of processing time.
+Once generated, these snapshots enable efficient verification of candidate patches, with each patch verification typically completing in seconds.
+Snapshot Restoration.
+At verification time, the appropriate snapshot is loaded and its contents restored into a temporary, isolated worktree that exactly replicates the post-build file structure of the original compilation.
+Files are fetched from the global CAS and written to disk.
+This process is stateless and fully parallelisable, enabling fast switching across historical commits without requiring repeated builds.
+Empirically, Eleanstic restores a full Lean environment in under one second on SSD-backed systems.
+To support long-term scalability, Eleanstic performs aggressive global deduplication across all snapshots.
+Across 3734 commits spanning August 2023 to March 2025, it compresses 15.6 TB of build artifacts into just 1.1 TB, yielding a storage reduction of over 93%.
+This compression enables persistent caching of every historical state required for evaluation, without exhausting disk resources.
+Patch Verification.
+After restoring the environment, the model-generated patch is applied to the target Lean file, and the resulting content is saved to a temporary file.
+Verification follows a strict file-level isolation policy: rooted in the restored Mathlib worktree, the Lean compiler compiles only the temporary file without affecting any files in the restored environment.
+Additionally, any temporary file that attempts to import the target file is discarded to prevent unintended leakage across file boundaries.
+A patch is considered syntactically valid if it compiles successfully within the corresponding version of Mathlib4 and does not emit warnings.
+Appendix C Additional Experimental Results and Analysis
+C.1 Model Performance Across pass@k Values
+Figure 10: Model performance across pass@k values. Left: Lean Compile pass@$k$ (syntactic validity). Right: LLM Judge pass@$k$ (semantic correctness conditioned on compilation).
+Figure 10 illustrates model behaviour across different pass@$k$ thresholds.
 The left panel reports syntactic success (Lean Compile), while the right panel shows semantic success (LLM-Judge) conditioned on compilation.
-o3-mini consistently outperforms other models across $k$ values in both dimensions, confirming its broad sampling diversity and task coverage.
-Claude Sonnet 3.7 (thinking) remains flat across $k$, reflecting its low-variance generation strategy.
-Figure 7: Model performance across pass@$k$ values.
-Left: Lean Compile pass@$k$ (syntactic validity); Right: LLM Judge pass@$k$ (semantic correctness conditioned on compilation).
-The gap between the two reflects semantic degradation.
-Notably, even the best-performing model—o3-mini—succeeds on only about one-third of the benchmark tasks, despite the fact that all tasks are constrained to single-file edits of no more than 100 lines.
-This highlights the substantial gap that remains between present capabilities and the demands of real-world proof engineering.
-In the following sections, we analyze this failure spectrum in greater detail by examining the structural and semantic robustness of model outputs.
-5.3 Structural Robustness and Task-Scale Sensitivity
-While overall success rates reveal substantial variation across models, they do not capture how model performance scales with the structural complexity of the underlying tasks.
-In this section, we analyze the extent to which models can handle edits of increasing size and structural scope, assess their ability to produce appropriately scaled patches, and evaluate their robustness across task types, difficulty levels, and mathematical domains.
-Performance Degradation on Larger Edits.
-We first group benchmark tasks by the size of their ground-truth patches and evaluate model pass@16 within each bin.
-As shown in Figure 8, most models exhibit sharp performance degradation as the size of the target edit increases.
-While all models achieve moderate to high success rates on small-scale edits (1–5 lines), performance drops significantly for medium (11–20 lines) and large-scale (>20 lines) tasks.
-In particular, o3-mini is the only model to retain non-trivial success on edits exceeding 50 lines, with a pass@16 about 25%, while all other models collapse to near-zero in this regime.
-These results highlight the inability of most LLMs to scale beyond localized modifications.
-Figure 8: Model pass@16 success rate grouped by ground-truth patch size.
-Larger edits cause sharp performance degradation for most models.
-Only o3-mini maintains partial success on tasks with more than 20 lines of changes.
-Patch Length Alignment with Ground-Truth.
-Figure 9 evaluates whether model-generated patches exhibit scale-aware behavior by comparing their length to the corresponding human edits.
-Most models over-edit small tasks (1–5 lines), producing longer-than-needed patches.
-As task size increases, generated patch sizes tend to align more closely with ground truth—except in models like GPT-4o and Doubao 1.5 Pro, which increasingly under-edit larger patches.
-This under-generation corresponds to their low semantic success on structurally complex tasks.
-In contrast, o3-mini and other strong models (Claude Sonnet, DeepSeek variants) exhibit broader variance and better coverage across patch sizes.
-Their outputs span both under- and over-editing, reflecting more diverse structural hypotheses.
-Notably, o3-mini frequently over-edits larger patches, which increases numeric deviation but improves semantic success by avoiding missing components—the dominant failure mode among conservative models.
-These results suggest that scale diversity, rather than strict length matching, is more critical for robust proof engineering performance.
+Gemini 2.5 Pro Preview achieves the highest semantic pass@$k$ across all values of $k$, reflecting consistent success in generating instruction-aligned patches.
+On the syntactic side, o3-mini leads at higher $k$ values due to more diverse sampling, though its semantic pass@$k$ grows more slowly in comparison.
+Claude Sonnet 3.7 (thinking) maintains stable but moderate gains in both dimensions.
+These trends highlight a persistent syntax–semantics gap: models that perform well in compilation do not always satisfy task intent.
+Sampling diversity improves raw compilation rates but does not necessarily translate to semantic alignment, underscoring the need for instruction-aware generation strategies.
+C.2 Patch Length Alignment with Ground-Truth
 Figure 11: Patch length alignment across models (log-log scale).
@@
 The dashed diagonal indicates ideal alignment.
-Stratified Performance by Task Type, Difficulty, and Domain.
-To further investigate structural sensitivity, we evaluate model performance across different task categories, difficulty levels, and mathematical domains.
-Figure 10 presents a unified heatmap showing evaluation pass@16 for each model across these dimensions.
-Tasks labeled as bug fixes are handled reasonably well by all models, likely due to their localized and low-risk nature.
-Refactor tasks reveal stronger model differentiation: Claude variants perform relatively well due to their structural preservation bias, whereas GPT-4o and Doubao underperform.
-Feature addition tasks are the most challenging, requiring new definitions, theorems, and often multi-declaration coherence; only o3-mini exhibits notable success in this category.
-Difficulty-based analysis further reveals a sharp capability cliff.
-While all models succeed on easy tasks, only o3-mini retains significant performance on medium and hard cases.
-Hard tasks, involving large edits or cross-declaration dependencies, remain virtually unsolved by all models except o3-mini.
-Finally, we observe domain-based variation: linear algebra and category theory consistently challenge most models, while tasks from algebra or data modules are comparatively tractable.
-These differences reflect varying structural depth and abstraction levels across formal math subfields.
-Figure 10: Heatmap of evaluation pass@16 success rates across different task types, difficulty levels, and mathematical modules.
-Column widths represent the proportion of each class in the test set.
-Color intensity indicates model performance.
-Summary.
-Taken together, these results indicate that most current models lack the structural robustness necessary for realistic proof engineering.
-Their success is largely limited to small-scale tasks and structurally shallow modifications.
-As patch size increases or task complexity deepens, performance degrades sharply.
-Even when a model produces a syntactically valid edit, it often fails to match the structural scope expected of a real engineer.
-In the next section, we examine whether these structural failures also correspond to deeper semantic errors.
-5.4 Semantic Robustness and Failure Characterization
-Despite producing syntactically valid Lean code, many model-generated patches fail to fulfill the semantic requirements expressed in the instruction.
-In this section, we examine the nature of these semantic failures in detail.
-We analyze patch repairability, categorize common failure modes, and compare error distributions across models to identify systematic weaknesses in current approaches.
-Structural Repairability of Model-Generated Patches.
-Raw patches generated by LLMs often contain formatting inconsistencies, invalid anchors, or malformed context that prevent successful application.
-To assess structural robustness, we measure patch application success before and after applying two forms of automated repair: exact match and fuzzy match (See Appendix A for full technical details.).
-As shown in Table 3, raw apply success rates remain low for all models, ranging from 12% to 16%, indicating poor structural precision in the initial output.
-Exact match repair offers moderate improvement for DeepSeek models, but fails to resolve structural errors in o3-mini or GPT-4o outputs.
-Fuzzy match repair significantly boosts success across the board, particularly for DeepSeek R1, Doubao, and Claude Sonnet 3.7.
-These results highlight the necessity of structural repair mechanisms in enabling meaningful evaluation and practical usability.
+Figure 11 evaluates whether model-generated patches exhibit scale-aware behaviour by comparing their length to the corresponding human edits.
+Most models still tend to over-edit small tasks (fewer than 8 lines of substantive changes, excluding blanks and comments), generating unnecessarily long patches.
+As task size increases, models like Gemini 2.5 Pro Preview and both Claude Sonnet and DeepSeek variants show strong alignment with the ground truth, maintaining near-diagonal trends and low mean squared error (MSE).
+In contrast, GPT-4o, o3-mini and Doubao 1.5 Pro significantly under-edit across the board, especially on larger edits, resulting in high patch-length MSE.
+This misalignment correlates with a higher frequency of quality defects and component omissions.
+These results suggest that while strict length matching is not a prerequisite for semantic correctness, extreme deviation—especially persistent over-editing—can lead to incomplete or misaligned edits.
+Models that strike a balance between scale-awareness and structural flexibility, such as Claude Sonnet, Gemini and DeepSeek V3, tend to perform more robustly in proof engineering tasks.
+C.3 DiffRepair: Patch Normalization System
+Model outputs are produced in unified diff format, which often contains structurally imprecise or misaligned edits (e.g., incorrect line anchors, out-of-date contexts).
+To ensure reliable application and verification, we introduce a post-processing phase called DiffRepair, which rewrites imprecise diffs into structurally consistent patches.
+This mechanism operates in two phases: first, it localises the intended edit position using fuzzy anchor matching; then it reconstructs the patch with preserved structural and contextual integrity.
+Intent Localisation.
+Given a noisy diff, DiffRepair first identifies where in the original Lean file the intended modification should take place.
+It uses a multi-granularity fuzzy matching strategy based on contextual anchors surrounding each diff hunk.
+Anchors are matched using exact or prefix-based line similarity, normalized token-level matching with whitespace tolerance, and structural proximity emphasising declaration-aligned contexts.
+Matches are scored using a dynamic programming model optimising for structural alignment and modification preservation.
+Patch Reconstruction.
+Once the target region is localised, DiffRepair reconstructs a clean diff by re-aligning added and deleted lines to structurally valid positions, augmenting missing context to satisfy unified diff constraints, resolving line number offsets and conflicting hunk overlaps, and merging multiple hunks into a minimal, coherent patch that preserves original semantics.
+This approach ensures that evaluation focuses on semantic correctness rather than superficial formatting issues.
 Model Raw Apply Success Exact Match Repair Fuzzy Match Repair
@@
 Table 6: Patch success rates before and after structural repair. Exact match applies conservative anchor correction; fuzzy match enables tolerant application using semantic alignment.
-Semantic Error Taxonomy and Cross-Model Distribution.
-To understand why many syntactically valid patches nonetheless fail semantically, we analyze failure cases using a taxonomy of five high-level semantic error types (Table 4).
-This categorization enables structured evaluation across models and highlights recurring patterns of semantic fragility.
-Error Class Description
-Component Omission Required definitions, proofs, or examples are missing
-Incompleteness Issue The edit is structurally partial, unfinished, or placeholder
-Mathematical Error The logic is invalid or inconsistent with Lean’s formal semantics
-Quality Defect Redundant or unabstracted edits, or lack of documentation
-Requirement Mismatch Misalignment with the intended instruction
-Table 4: Semantic error classes and their typical manifestations.
-Figure 11: Distribution of semantic error types across models (log-scaled y-axis).
-Component omission is the most common error overall.
-GPT-4o and o3-mini frequently produce incomplete or mathematically incorrect edits, while Claude Sonnet 3.7 (thinking) exhibits the lowest error incidence.
-Figure 11 presents the frequency of each error type as a proportion of failure cases per model.
-The most pervasive issue is component omission, accounting for a significant share of failures across all models.
-This indicates a widespread tendency to under-produce or neglect structurally necessary elements—such as required definitions, lemmas, or proofs—even when the instruction clearly specifies them.
-Beyond omission, GPT-4o and o3-mini suffer from notably high rates of incompleteness and mathematical errors.
-These failures suggest limited planning ability when tasked with multi-step edits, and difficulty in maintaining formal soundness under structural pressure.
-By contrast, Claude Sonnet 3.7 (thinking) demonstrates a markedly lower error profile across all categories.
-Summary.
-Our findings indicate that semantic failures in LLM-generated patches are not random but structurally patterned.
-The majority stem from incomplete or underspecified edits, often due to scale underestimation or partial instruction comprehension.
-While fuzzy repair can address syntactic alignment, it cannot compensate for missing logic or misaligned intent.
-Addressing these issues will require models that plan multi-step edits, verify task satisfaction, and iteratively refine their output—an agenda we return to in our discussion of future directions.
-6 Conclusion
-This work introduces the paradigm of Automated Proof Engineering (APE) and presents APE-Bench I, the first large-scale benchmark grounded in real-world proof development workflows from Mathlib4.
-By reframing formal mathematics as instruction-driven code editing, and combining Lean compilation with LLM-based semantic evaluation, APE-Bench I offers a rigorous and scalable framework for evaluating the structural and semantic capabilities of LLMs in realistic proof engineering tasks.
-Our empirical results reveal significant performance gaps among state-of-the-art models, particularly under structurally complex scenarios, and underscore the need for future systems that can plan, verify, and refine edits across evolving formal libraries.
-This benchmark establishes a strong foundation for developing next-generation autonomous agents capable of practical and scalable formal proof maintenance.
-Appendix A DiffRepair: Fault-Tolerant Patch Recovery for LLM Outputs
-A.1 Motivation and Overview
-Model-generated patches in unified diff format frequently fail to apply cleanly to real-world codebases due to misaligned line anchors, outdated contextual information, or excessive reliance on idealized formatting.
-This undermines the practical viability of LLM-generated edits in structured formal environments like Lean.
-To address this, we develop DiffRepair, a two-stage recovery system that transforms noisy model diffs into clean, structure-preserving, and context-valid patches that can be compiled and semantically evaluated.
-DiffRepair adopts a tolerant matching and reconstruction strategy.
-Rather than enforcing exact line-based alignment, it reconstructs the intent behind the edit, anchors it in the correct structural region, and regenerates a diff conforming to unified format constraints.
-This allows it to serve as a bridge between neural code generation and downstream compilation pipelines.
-A.2 Algorithmic Workflow
-1. Intent Localization.
-Given a noisy diff, DiffRepair first identifies where in the original Lean file the intended modification should take place.
-It uses a multi-granularity fuzzy matching strategy based on contextual anchors surrounding each diff hunk.
-Anchors are matched using:
-  • Exact or prefix-based line similarity.
-  • Normalized token-level matching with whitespace and casing tolerance.
-  • Structural proximity, emphasizing declaration-aligned contexts (e.g., def, theorem blocks).
-Matches are scored using a dynamic programming model optimizing for structural alignment and modification preservation, ensuring minimal distortion to the intended patch intent.
-2. Patch Reconstruction.
-Once the target region is localized, DiffRepair reconstructs a clean diff by:
-  • Re-aligning added and deleted lines to structurally valid positions.
-  • Augmenting missing context to satisfy unified diff constraints.
-  • Resolving line number offsets and conflicting hunk overlaps.
-  • Merging multiple hunks into a minimal, coherent patch that preserves original semantics.
-Reconstructed patches are guaranteed to apply via standard tools such as patch or git apply, enabling seamless integration into downstream pipelines.
-A.3 Design Features
-DiffRepair is designed to satisfy the following desiderata:
-  • Robustness to noisy generation: Handles line number inaccuracies, formatting inconsistencies, and incomplete context.
-  • Structure preservation: Ensures generated diffs align with Lean’s structural boundaries and avoid cross-declaration contamination.
-  • Automation-ready: Requires no human-in-the-loop intervention and supports batch repair at scale.
-  • Efficiency: Performs fast multi-hunk matching and reconstruction with low overhead.
+As shown in Table 6, raw apply success rates remain low for all models, ranging from 12% to 16%, indicating poor structural precision in the initial output.
+Exact match repair offers moderate improvement for DeepSeek models, but fails to resolve structural errors in o3-mini or GPT-4o outputs.
+Fuzzy match repair significantly boosts success across the board, particularly for DeepSeek R1, Doubao, and Claude Sonnet 3.7.
+These results highlight the necessity of structural repair mechanisms for enabling meaningful evaluation and practical usability.
+Appendix D Illustrative Examples
+Source code:
+Copyright (c) 2017 Mario Carneiro. All rights reserved.
+Released under Apache 2.0 license as described in the file LICENSE.
+Authors: Mario Carneiro
+import Mathlib.Data.Fintype.EquivFin
+import Mathlib.Data.Finset.Option
+# fintype instances for option
+assert_not_exists MonoidWithZero MulAction
+open Function
+open Nat
+universe u v
+variable β : Type*}
+open Finset Function
+instance : Type*} [Fintype α] : Fintype (Option α) :=
+Finset.insertNone univ, fun a => by simp
+theorem univ_option : Type*) [Fintype α] : (univ : Finset (Option α)) = insertNone univ :=
+rfl
+@[simp]
+theorem Fintype.card_option : Type*} [Fintype α] :
+Fintype.card (Option α) = Fintype.card α + 1 :=
+(Finset.card_cons (by simp)).trans <| congr_arg _ (card_map _) rfl
+/– If Option α‘ is a Fintype then so is ‘α‘ -/
+def fintypeOfOption : Type*} [Fintype (Option α)] : Fintype α :=
+Finset.eraseNone (Fintype.elems := Option α)), fun x =>
+mem_eraseNone.mpr (Fintype.complete (some x))⟩
+/– A type is a Fintype if its successor (using Option‘) is a Fintype‘. -/
+def fintypeOfOptionEquiv [Fintype α] (f : α Option β) : Fintype β :=
+haveI := Fintype.ofEquiv _ f
+fintypeOfOption
+namespace Fintype
+/– A recursor principle for finite types, analogous to Nat.rec‘. It effectively says
+that every Fintype is either Empty or Option α‘, up to an Equiv‘. -/
+def truncRecEmptyOption {P : Type u Sort v} (of_equiv : β}, α β P α P β)
+(h_empty : P PEmpty) (h_option : {α} [Fintype α] [DecidableEq α], P α P (Option α))
+: Type u) [Fintype α] [DecidableEq α] : Trunc (P α) := by
+suffices n : ℕ, Trunc (P (ULift <| Fin n)) by
+apply Trunc.bind (this (Fintype.card α))
+intro h
+apply Trunc.map _ (Fintype.truncEquivFin α)
+intro e
+exact of_equiv (Equiv.ulift.trans e.symm) h
+apply ind where
+Porting note: do a manual recursion, instead of induction tactic,
+to ensure the result is computable
+/– Internal induction hypothesis -/
+ind : n : ℕ, Trunc (P (ULift <| Fin n))
+| Nat.zero => by
+have : card PEmpty = card (ULift (Fin 0)) := by simp only [card_fin, card_pempty,
+card_ulift]
+apply Trunc.bind (truncEquivOfCardEq this)
+intro e
+apply Trunc.mk
+exact of_equiv e h_empty
+| Nat.succ n => by
+have : card (Option (ULift (Fin n))) = card (ULift (Fin n.succ)) := by
+simp only [card_fin, card_option, card_ulift]
+apply Trunc.bind (truncEquivOfCardEq this)
+intro e
+apply Trunc.map _ (ind n)
+intro ih
+exact of_equiv e (h_option ih)
+Porting note: due to instance inference issues in SetTheory.Cardinal.Basic
+I had to explicitly name h_fintype in order to access it manually.
+was ‘[Fintype α]‘
+/– An induction principle for finite types, analogous to Nat.rec‘. It effectively says
+that every Fintype is either Empty or Option α‘, up to an Equiv‘. -/
+@[elab_as_elim]
+theorem induction_empty_option {P : : Type u) [Fintype α], Prop}
+(of_equiv : β) [Fintype β] (e : α β), @P α (@Fintype.ofEquiv α β < _ > e.symm) @P β )
+(h_empty : P PEmpty) (h_option : (α) [Fintype α], P α P (Option α)) : Type u)
+[h_fintype : Fintype α] : P α := by
+obtain p :=
+let f_empty := fun i => by convert h_empty
+let h_option : : Type u} [Fintype α] [DecidableEq α],
+(∀ (h : Fintype α), P α) (h : Fintype (Option α)), P (Option α) := by
+rintro α hα - Pα hα’
+convert h_option α (Pα _)
+@truncRecEmptyOption (fun α => h, @P α h) (@fun α β e hα hβ => @of_equiv α β hβ e (hα _))
+f_empty h_option α _ (Classical.decEq α)
+exact p _
+end Fintype
+/– An induction principle for finite types, analogous to Nat.rec‘. It effectively says
+that every Fintype is either Empty or Option α‘, up to an Equiv‘. -/
+theorem Finite.induction_empty_option {P : Type u Prop} (of_equiv : β}, α β P α P β)
+(h_empty : P PEmpty) (h_option : {α} [Fintype α], P α P (Option α)) : Type u)
+[Finite α] : P α := by
+cases nonempty_fintype α
+refine Fintype.induction_empty_option ?_ ?_ ?_ α
+exacts [fun α β _ => of_equiv, h_empty, @h_option]
+Instructions:
+- Task 1: Modernizing a Recursion Principle for Finite Types
+Consider the recursion principle for finite types that states: "To define a function or prove a property for all finite types, it suffices to handle the empty set case, show how to extend from a finite set to that set plus one extra element, and respect type equivalences."
+This principle is implemented using induction on the cardinality of finite types.
+Refactor the implementation by replacing the manually defined recursive function with Leans standard induction tactic while preserving the mathematical content.
+Specifically:
+1. Remove the manual recursion approach that uses the where syntax and explicit pattern matching.
+2. Introduce the natural number parameter and apply induction directly.
+3. In the inductive step, use the induction hypothesis provided by the tactic rather than making an explicit recursive call.
+4. Ensure that all mathematical reasoning about cardinality equivalences remains intact.
+The refactoring should modernize the code while maintaining the same mathematical structure and reasoning.
+Generation Response of DeepSeek-R1:
+<think>
+Okay, I need to tackle this task of refactoring the recursion principle for finite types in the given Lean4 code.
+The main goal is to replace the manual recursion approach with Leans standard induction tactic while preserving the mathematical content.
+Let me try to break this down step by step.
+First, I need to understand the existing code.
+The truncRecEmptyOption function uses a manual recursion approach with ind defined using the where syntax and explicit pattern matching.
+The comments mention that this was done to ensure computability, but the task now wants to switch to using Leans induction tactic.
+The key points from the task are:
+1. Remove manual recursion (‘where syntax and explicit pattern matching).
+2. Introduce a natural number parameter and apply induction directly.
+3. Use the induction hypothesis provided by the tactic instead of explicit recursive calls.
+4. Keep the mathematical reasoning intact, especially around cardinality equivalences.
+So, looking at the truncRecEmptyOption function, the manual recursion is handled by the ind function inside the apply ind where block. The ind function is defined with Nat.zero and Nat.succ n cases.
+To refactor this, I need to replace this manual recursion with a standard induction tactic.
+That means structuring the proof so that Leans induction tactic can handle the recursion based on the natural number parameter representing the cardinality.
+The suffices block is setting up the proof by showing that for all natural numbers n‘, the property P holds for ULift (Fin n)‘.
+Then, it binds the result of this to the equivalence between ‘α‘ and Fin (Fintype.card α)‘.
+The apply ind where part is the manual recursion that needs to be replaced. Instead, I should introduce a natural number parameter (probably the cardinality) and use induction on that.
+In the Nat.zero case, the current code uses truncEquivOfCardEq to get an equivalence between PEmpty and ULift (Fin 0)‘, then applies of_equiv with h_empty‘.
+Similarly, in the Nat.succ n case, it shows the equivalence between Option (ULift (Fin n))‘ and ULift (Fin n.succ)‘ and uses h_option with the induction hypothesis from ind n‘.
+So, to use the induction tactic, Ill need to structure the proof such that:
+- The induction is on the natural number n (cardinality).
+- In the base case (‘n = 0‘), handle the empty set scenario.
+- In the inductive step (‘n.succ‘), use the induction hypothesis to build the proof for n.succ‘.
+The induction tactic should provide the induction hypothesis automatically, so I dont need to make explicit recursive calls.
+I also need to ensure that all the equivalences and mathematical reasoning about cardinality are preserved. The truncEquivOfCardEq and Equiv.ulift are crucial here for maintaining those properties.
+Let me sketch out the changes:
+1. Remove the apply ind where block and the ind function.
+2. Introduce a natural number parameter for induction, likely using Fintype.card α‘.
+3. Use induction on this natural number.
+4. In each case, use the existing logic with truncEquivOfCardEq and Equiv.ulift to handle the equivalences.
+5. Ensure that the h_empty and h_option cases are properly applied with the induction hypothesis.
+Possible challenges:
+- Ensuring that the induction hypothesis is correctly applied in the inductive step.
+- Maintaining the correct use of truncEquivOfCardEq and Equiv.ulift when restructuring the code.
+- Making sure that all the required instances and typeclass parameters are correctly inferred or provided.
+Testing the changes would involve checking that the refactored code compiles and that the mathematical content remains unchanged.
+Since the task specifies preserving the mathematical structure, its crucial that the equivalences and cardinality reasoning are not altered.
+Another thing to note is the @[elab_as_elim]‘ attribute on induction_empty_option‘, which suggests that this theorem is intended to be used as an elimination principle.
+The refactoring should ensure that this attribute is still appropriate and that the theorems usability remains intact.
+In summary, the main steps are:
+- Replace manual recursion with induction tactic.
+- Structure the proof around a natural number parameter for induction.
+- Preserve all equivalences and cardinality reasoning.
+- Ensure that the induction hypothesis is correctly utilized in the inductive step.
+By carefully following these steps and checking against the original codes logic, I should be able to produce the required unified diff patch.
+</think>
+To address the task requirements, well refactor the truncRecEmptyOption function by replacing manual recursion with Leans induction tactic while preserving the mathematical content.
+Heres the step-by-step explanation and the corresponding diff:
+**Step-by-Step Explanation:**
+1. **Introduce Natural Number Parameter:**
+- The original code manually recurses on the cardinality of the finite type. Well instead use a natural number parameter n and perform induction on it.
+2. **Apply Induction Tactic:**
+- Replace the manual ind function with induction n to leverage Leans built-in induction tactic. This simplifies the code and uses the induction hypothesis provided by the tactic.
+3. **Base Case (‘n = 0‘):**
+- Handle the empty set case by using truncEquivOfCardEq to establish the equivalence between PEmpty and ULift (Fin 0)‘, then apply h_empty via of_equiv‘.
+4. **Inductive Step (‘n + 1‘):**
+- For the successor case, use truncEquivOfCardEq to equate Option (ULift (Fin n))‘ with ULift (Fin n.succ)‘.
+- Apply the induction hypothesis ih (from the induction tactic) and h_option to build the proof for n + 1‘.
+5. **Preserve Equivalences:**
+- Maintain the use of Equiv.ulift and truncEquivOfCardEq to ensure the cardinality reasoning remains intact.
+6. **Cleanup and Simplification:**
+- Remove the now-unnecessary ind function and associated comments.
+- Adjust variable bindings and scopes to match the new structure.
+**Unified Diff Patch:**
+“‘diff
+diff git a/Mathlib/Data/Fintype/Option.lean b/Mathlib/Data/Fintype/Option.lean
+index 1234567..abcdefg 100644
+a/Mathlib/Data/Fintype/Option.lean
+$+$ b/Mathlib/Data/Fintype/Option.lean
+@@ -52,29 +52,23 @@ def truncRecEmptyOption {P : Type u Sort v} (of_equiv : β}, α
+(h_empty : P PEmpty) (h_option : {α} [Fintype α] [DecidableEq α], P α P (Option α))
+: Type u) [Fintype α] [DecidableEq α] : Trunc (P α) := by
+suffices n : ℕ, Trunc (P (ULift <| Fin n)) by
+- apply Trunc.bind (this (Fintype.card α))
+- intro h
+- apply Trunc.map _ (Fintype.truncEquivFin α)
+- intro e
+- exact of_equiv (Equiv.ulift.trans e.symm) h
+- apply ind where
+- Porting note: do a manual recursion, instead of induction tactic,
+- to ensure the result is computable
+- /– Internal induction hypothesis -/
+- ind : n : ℕ, Trunc (P (ULift <| Fin n))
+- | Nat.zero => by
+- have : card PEmpty = card (ULift (Fin 0)) := by simp only [card_fin, card_pempty,
+- card_ulift]
+- apply Trunc.bind (truncEquivOfCardEq this)
+- intro e
+- apply Trunc.mk
+- exact of_equiv e h_empty
+- | Nat.succ n => by
+- have : card (Option (ULift (Fin n))) = card (ULift (Fin n.succ)) := by
+- simp only [card_fin, card_option, card_ulift]
+- apply Trunc.bind (truncEquivOfCardEq this)
+- intro e
+- apply Trunc.map _ (ind n)
+- intro ih
+- exact of_equiv e (h_option ih)
++ exact Trunc.bind (this (Fintype.card α)) fun h =>
++ Trunc.map (fun e => of_equiv (Equiv.ulift.trans e.symm) h) (Fintype.truncEquivFin α)
++ intro n
++ induction n with
++ | zero =>
++ have : card PEmpty = card (ULift (Fin 0)) := by simp only [card_fin, card_pempty, card_ulift]
++ apply Trunc.bind (truncEquivOfCardEq this)
++ intro e
++ apply Trunc.mk
++ exact of_equiv e h_empty
++ | succ n ih =>
++ have : card (Option (ULift (Fin n))) = card (ULift (Fin n.succ)) := by
++ simp only [card_fin, card_option, card_ulift]
++ apply Trunc.bind (truncEquivOfCardEq this)
++ intro e
++ apply Trunc.map _ ih
++ intro ih
++ exact of_equiv e (h_option ih)
+Porting note: due to instance inference issues in SetTheory.Cardinal.Basic
+I had to explicitly name h_fintype in order to access it manually.
+**Explanation of the Diff:**
+- **Removed Manual Recursion (‘ind function):**
+- The original codes apply ind where block and manually defined ind function are deleted.
+- **Introduced induction Tactic:**
+- The induction n with block replaces manual recursion, using Leans built-in induction.
+- **Base Case (‘zero‘):**
+- Handles n = 0‘ by establishing equivalence between PEmpty and ULift (Fin 0)‘, then applying h_empty‘.
+- **Inductive Step (‘succ n ih‘):**
+- Uses ih (induction hypothesis) provided by the induction tactic.
+- Establishes equivalence between Option (ULift (Fin n))‘ and ULift (Fin n.succ)‘.
+- Applies h_option with ih to build the step case.
+- **Preserved Mathematical Content:**
+- All cardinality equivalence checks via truncEquivOfCardEq and Equiv.ulift remain intact.
+- The structure of the proof aligns with the original mathematical reasoning.
+This refactoring modernizes the code by leveraging Leans induction tactic, improving readability and maintainability while preserving the original mathematical structure.
+Judgement Response of Claude Sonnet 3.7 (thinking):
+<think>
+Lets analyze the proposed implementation to determine if it fulfills the task requirements. First, lets understand the original code and then evaluate the changes.
+## Original Code Understanding
+The original code defines a recursion principle for finite types called truncRecEmptyOption‘. This function takes:
+1. A way to transport through equivalences (‘of_equiv‘)
+2. A base case for the empty set (‘h_empty‘)
+3. An inductive step for adding one element via Option (‘h_option‘)
+4. A type ‘α‘ with Fintype and DecidableEq instances
+The implementation works by:
+1. Reducing the problem to showing the principle for Fin n types (standard finite types of size n)
+2. Using a manually defined recursive function ind to prove this for all natural numbers
+3. For the base case (n=0), it establishes an equivalence with the empty type
+4. For the inductive case, it uses the recursive hypothesis and establishes an equivalence between Option (Fin n)‘ and Fin (n+1)‘
+The manual recursion is done using the where clause with explicit pattern matching on natural numbers.
+## Task Requirements
+The task asks to modernize this recursion principle by:
+1. Removing the manual recursion approach with where syntax
+2. Using Leans standard induction tactic
+3. Using the induction hypothesis provided by the tactic
+4. Preserving all mathematical reasoning
+## Changes in the Proposed Implementation
+Lets examine the changes:
+1. **Removed manual recursion**: The implementation removed the ind function defined with the where clause and explicit pattern matching.
+2. **Introduced induction tactic**: The implementation now uses induction n with to directly perform induction on the natural number parameter.
+3. **Using induction hypothesis**: Instead of making an explicit recursive call via ind n‘, the implementation now uses the induction hypothesis ih provided by the tactic.
+4. **Mathematical reasoning preserved**: The mathematical content about cardinality equivalences remains intact. The same logic is used in both the base case and the inductive step:
+- In the base case, we establish an equivalence between PEmpty and ULift (Fin 0)‘
+- In the inductive step, we establish an equivalence between Option (ULift (Fin n))‘ and ULift (Fin n.succ)‘
+5. **Code cleanup**: The implementation also simplifies some nested Trunc operations by using more concise expressions:
+“‘lean
+exact Trunc.bind (this (Fintype.card α)) fun h =>
+Trunc.map (fun e => of_equiv (Equiv.ulift.trans e.symm) h) (Fintype.truncEquivFin α)
+This replaces the more verbose:
+“‘lean
+apply Trunc.bind (this (Fintype.card α))
+intro h
+apply Trunc.map _ (Fintype.truncEquivFin α)
+intro e
+exact of_equiv (Equiv.ulift.trans e.symm) h
+Overall, the changes maintain the same mathematical structure and reasoning while modernizing the approach to use Leans built-in induction tactic rather than a manually defined recursive function.
+## Evaluation
+Lets assess whether the implementation correctly fulfills the task requirements:
+1. **Removal of manual recursion**: The manual recursion using where and explicit pattern matching has been removed.
+2. **Introduction of induction tactic**: The implementation now uses induction n with to perform induction on the natural number parameter.
+3. **Use of induction hypothesis**: The implementation now uses the induction hypothesis ih provided by the tactic rather than making an explicit recursive call.
+4. **Preservation of mathematical reasoning**: All mathematical reasoning about cardinality equivalences remains intact. The same logic is used in both the base case and the inductive step.
+The implementation properly fulfills all the task requirements.
+The code has been modernized while maintaining the same mathematical structure and reasoning.
+The use of Leans induction tactic makes the code cleaner and more in line with modern Lean practices.
+</think>
+# Comprehensive Analysis of Lean4 Code Patch
+## Task Understanding
+The task requires modernizing a recursion principle for finite types by replacing a manual recursion approach with Leans standard induction tactic, while preserving the mathematical content.
+The goal is to update the code style while maintaining the same mathematical reasoning.
+## Original Code Analysis
+The original code implements truncRecEmptyOption‘, a recursion principle for finite types.
+It uses a manual recursive function definition with where syntax and explicit pattern matching on natural numbers:
+“‘lean
+apply ind where
+ind : n : ℕ, Trunc (P (ULift <| Fin n))
+| Nat.zero =>
+| Nat.succ n =>
+This function establishes that to define a property for all finite types, its sufficient to:
+1. Handle the empty set case
+2. Show how to extend from a finite set to that set plus one element
+3. Respect type equivalences
+The approach works by induction on the cardinality of the finite type, using equivalences between PEmpty and Fin 0‘, and between Option (Fin n)‘ and Fin (n+1)‘.
+## Patch Examination
+The patch makes the following changes:
+1. Removes the manual recursion using where syntax and explicit pattern matching
+2. Introduces direct parameter n and applies the induction tactic with induction n with
+3. Uses the induction hypothesis ih provided by the tactic rather than making explicit recursive calls
+4. Preserves all mathematical reasoning about cardinality equivalences
+5. Streamlines some code by combining nested Trunc operations into more concise expressions
+For example, this verbose nested block:
+“‘lean
+apply Trunc.bind (this (Fintype.card α))
+intro h
+apply Trunc.map _ (Fintype.truncEquivFin α)
+intro e
+exact of_equiv (Equiv.ulift.trans e.symm) h
+Is replaced with a more concise:
+“‘lean
+exact Trunc.bind (this (Fintype.card α)) fun h =>
+Trunc.map (fun e => of_equiv (Equiv.ulift.trans e.symm) h) (Fintype.truncEquivFin α)
+## Requirement Fulfillment Analysis
+The implementation successfully meets all task requirements:
+1. The manual recursion approach using where syntax has been eliminated
+2. The natural number parameter is introduced directly and induction is applied with the standard tactic
+3. The inductive step uses the tactic-provided induction hypothesis (‘ih‘) rather than an explicit recursive call
+4. All mathematical reasoning about cardinality equivalences remains intact
+## Implementation Quality Analysis
+The modernized implementation is:
+- More idiomatic to contemporary Lean style by using built-in tactics
+- Clearer in structure, as the induction flow is more explicit
+- More maintainable, as it follows standard patterns that Lean developers will be familiar with
+- Functionally equivalent to the original, preserving all mathematical content
+- Slightly more concise due to streamlined expression of Trunc operations
+## Overall Judgement
+“‘json
+"TaskEvaluations": {
+"Task 1": "Excellent"
+"FinalOverallGrade": "Excellent"